### v1

# Quantum space-time Poincaré inequality for Lindblad dynamics

Abstract.
We investigate the mixing properties of primitive hypocoercive Lindblad dynamics. By extending the variational framework originally developed for underdamped Langevin dynamics, we derive fully explicit and constructive exponential decay estimates for the convergence in the noncommutative $L^{2}$-norm. Our analysis relies on establishing a quantum analog of space-time Poincaré inequalities. To complement these hypocoercive estimates, we also analyze the limiting behavior of the spectral gap for Lindblad dynamics with a large coherent contribution, providing sharper convergence rate estimates in this asymptotic regime. A number of concrete examples are provided as applications of our theoretical results.
1. Introduction
Quantum systems inherently interact with their environments, leading to dissipative processes and decoherence. In the regime of weak system-environment coupling, these open quantum dynamics become Markovian and can be modeled through the Lindblad master equation, also known as quantum Markov semigroups (QMS) [breuer2002theory]. The Lindblad dynamics share a structural similarity with classical underdamped Langevin equations, involving Hamiltonian evolution alongside the fluctuation-dissipation terms that account for the environmental interactions (see Remark 1 for more details).
The Markovian assumption simplifies the mathematical treatment of open quantum dynamics and facilitates efficient numerical simulations. Consequently, the Lindblad equation is a widely applicable tool in various fields, including condensed matter physics [zhang2024driven, barthel2022superoperator], quantum information theory [kastoryano2013quantum, temme2010chi], and quantum computation [verstraete2009quantum, kastoryano2011dissipative]. With the rapid advancement of quantum computers and quantum algorithms in recent years, there has been a growing interest in applying Lindblad dynamics for quantum Gibbs sampling [chen2023quantum, chen2023efficient, ding2024efficient], ground state preparation [ding2023single], and solving nonconvex optimization problems [li2023quantum]. To analyze the efficiency of these Lindblad dynamics-based quantum algorithms, it is crucial to understand the mixing properties of the underlying dynamics, which is also of fundamental mathematical importance. Notably, it has been observed in [ding2023single] that adding a coherent term to the Lindblad dynamics, which breaks the detailed balance conditions (DBC), can often accelerate the convergence towards equilibrium. The primary goal of this work is to provide an explicit $L^{2}$-decay rate estimate for the primitive hypocoercive Lindblad dynamics, motivated by the variational methods [albritton2019variational, cao2023explicit] recently developed for kinetic Fokker-Planck equations.
1.1. Related works
The mixing time of a Markovian system is the duration required for the system to reach its stationary distribution from an arbitrary initial state. One of the most common and important approaches to bounding the mixing time is estimating the spectral gap of the Markov semigroup using $\chi^{2}$-divergence [diaconis1991geometric, temme2010chi]. It is well known that a system-size-independent spectral gap estimate only yields a $L^{1}$-mixing time of $t_{\rm mix}=\mathcal{O}(n)$, where $N=2^{N}$ is the dimension of the quantum state space. In order to achieve the rapid mixing $t_{\rm mix}=\mathcal{O}(\log(n))$, one must utilize the modified log-Sobolev inequality (MLSI), which has been explored in [kastoryano2013quantum]. Since then, substantial efforts have been made to estimate the spectral gap and MLSI constants for Lindblad dynamics, particularly those related to quantum spin systems. Kastoryano and Brandão [kastoryano2016quantum] considered the Gibbs state preparation of spin Hamiltonians with commuting local terms and demonstrated that the spectral gap of Davies generators is independent of system size if and only if the Gibbs state satisfies a strong clustering of correlations. Capel et al. [bardet2021modified, cuevas2019quantum, capel2020modified] investigated the MLSI for quantum spin lattice systems through quasi-factorization or approximate tensorization of the quantum relative entropy. In recent work [kochanowski2024rapid], rapid mixing was proved for a large class of geometrically-2-local Davies generators with commuting Hamiltonians. For the generic positivity of quantum MLSI, the notion of Ricci curvature lower bounds (geodesic convexity) has played an important role. Such a concept is pioneered by Carlen and Maas [carlen2014analog, carlen2017gradient], where they showed that a primitive QMS satisfying $\sigma$-GNS detailed balance condition is a gradient flow of the relative entropy with respect to some quantum Wasserstein distance. This allows us to consider the Ricci curvature of a QMS and its connection with MLSI, which was explored by Datta and Rouzé [datta2020relating]. The extension of the gradient flow structure and Ricci curvature notion to a general family of quantum convex relative entropies was investigated in [li2023interpolation], which gives the so-called quantum $p$-Beckner inequality [adamczak2022modified].
Quantum non-primitive models. The results mentioned above on the long-time behaviors of Lindblad dynamics are restricted to primitive QMS with detailed balance conditions. Extending these results to the non-primitive case is closely related to the concept of environment-induced decoherence [blanchard2003decoherence] (noting that for classical Markov processes, non-primitivity implies a degenerate diffusive generator). Bardet [bardet2017estimating] initiated the study of mixing (decoherence) properties for non-primitive Lindblad dynamics from the spectral gap (equivalently, the non-primitive Poincaré inequality), which has emerged as a very active research area in the past few years. In this setup, Gao et al. [gao2020fisher] introduced a new tensor-stable MLSI: the complete modified log-Sobolev inequality (CMLSI), proving its validity through the monotonicity of Fisher information. More recently, Gao and Rouzé [gao2021complete] showed that CMLSI holds for any finite-dimensional non-primitive GNS-detailed balanced QMS through a two-sided estimate for the relative entropy. Connections between CMLSI, entropic or geometric Ricci curvature, and gradient estimates have been explored in [li2020graph, wirth2021complete, brannan2022complete, wirth2021curvature]. We mention that the recent work [gao2022complete] provided a generic asymptotically tight lower bound for the CMLSI constant using the inverse of completely bounded return time and an improved data processing inequality.
Acceleration by irreversible Markov processes. As basic models in non-equilibrium statistical physics, the irreversible Markov dynamics has been studied extensively in the past few decades [hairer2009hot, eckmann2000non]. Recent findings indicate that breaking reversibility (detailed balance) can often accelerate the mixing of Markov processes, which is a favorable property for Monte Carlo Markov chain method for probability sampling. Various criteria were proposed to quantify the ergodic properties of the Markov process and compare the efficiency of the corresponding sampling method [fill1991eigenvalue, chatterjee2023spectral, hwang2005accelerating, peskun1973optimum, neal2004improving, rey2015variance, rey2015irreversible, ferre2020large]. Hwang et al. [hwang2005accelerating] proved that under some mild conditions, the spectral gap of a reversible diffusion can strictly decrease by adding a divergence-free linear drift, thereby accelerating convergence to equilibrium. Further analysis by Franke et al. [franke2010behavior] explored the limiting behavior of the spectral gap of the Laplace-Beltrami operator with a growing linear drift vector field on a compact Riemannian manifold. See also [fill1991eigenvalue, chatterjee2023spectral] for the spectral gap analysis for finite-state Markov chains. While the spectral properties of a Markov generator provide a useful measure of ergodicity, from a Monte Carlo sampling perspective, it is more informative to consider the statistics of the ergodic average $t^{-1}\int_{0}^{t}f(X_{s})\,\mathrm{d}s$ or the empirical measure $t^{-1}\int_{0}^{t}\delta_{X_{s}}\,\mathrm{d}s$. For example, Dupuis et al. [dupuis2012infinite] utilized the large deviation rate function to compare the performance of Markov samplers and the convergence of ergodic averages for parallel tempering-type algorithms. Later, Rey-Bellet and Spiliopoulos [rey2015irreversible, rey2015variance] considered irreversible Langevin samplers and demonstrated that adding a growing drift decreases the asymptotic variance of the estimator and increases the large deviations rate function of the empirical measure. We refer the readers to [ferre2020large] and references therein for more recent results.
Hypocoercivity theory. Another framework for analyzing the exponentially fast convergence of irreversible hypoelliptic Markov processes is the hypocoercivity theory, which addresses models with a degenerate diffusion (elliptic) component, such as the kinetic Fokker-Planck equations (Remark 1). The foundational idea, initiated by Hérau [desvillettes2001trend, herau2004isotropic, herau2006hypocoercivity] and then systematically developed by Villani [villani2009hypocoercivity], is to split the generator into reversible and irreversible parts and then quantify they interact with each other. Due to the degenerate ellipticity, the system is not coercive in the standard $L^{2}$-norm. In order to carry the dissipation from the momentum direction to whole space, it is often necessary to twist the reference norm. The $H^{1}$-hypocoercivity has been considered in [villani2009hypocoercivity, mouhot2006quantitative], directly implying convergence in the $L^{2}$-norm. However, the resulting rate is implicit and tends to be suboptimal. A direct approach for the $L^{2}$-convergence was investigated by Dolbeault, Mouhot and Schmeiser (DMS) [dolbeault2015hypocoercivity, dolbeault2009hypocoercivity], based on a modified $L^{2}$-norm (i.e., Lyapunov function). Its underlying idea was inspired from [herau2006hypocoercivity] and revisited in [grothaus2016hilbert]. A series of works [arnold2014sharp, achleitner2015large, arnold2020propagator] by Arnold et al. constructed another modified norm and derived the optimal convergence rates, but it relies on algebraic properties and the spectral information of the involved diffusion operator.
A new variational hypocoercivity framework was recently proposed by Armstrong and Mourrat [albritton2019variational] by using the space-time adapted Poincaré inequalities, without changing the $L^{2}$-scalar product. It was later extended by Cao et al. [cao2023explicit] for underdamped Langevin dynamics with a confining potential, and by Lu and Wang [lu2022explicit] for piecewise deterministic Markov processes. Compared to DMS entropy functional approach [dolbeault2015hypocoercivity, dolbeault2009hypocoercivity], the variational framework directly analyzes the dynamics and can potentially provide optimal estimates on the convergence rate. For instance, in the case of underdamped Langevin dynamics, assuming the Hessian of the potential is bounded from below and choosing the friction coefficient $\gamma=\mathcal{O}(\sqrt{m})$, [cao2023explicit] obtained the optimal rate $\mathcal{O}(\sqrt{m})$ as $m\to 0$, whereas the DMS approach only yields a rate $\mathcal{O}(m^{5/2})$ [eberle2024non]*Remark 29, where $m$ is the Poincaré constant for the target distribution. For a more comprehensive literature review on hypocoercivity, see [brigati2023construct, bernard2022hypocoercivity, cao2023explicit]. It is also worth mentioning that this variational approach with space-time Poincaré inequalities was recently rephrased and simplified by Eberle and Lörler using a second-order lift framework [eberle2024non].
1.2. Main results and discussions
As reviewed above, while significant progress has been made in understanding classical irreversible dynamics and hypocoercivity theory, most results concerning the convergence of quantum Markov processes assume detailed balance conditions (DBC). Results for quantum irreversible dynamics are quite limited. Temme et al. [temme2010chi] considered the convergence of the quantum channel with or without DBC by singular value analysis. More recently, Fang et al. [fang2024mixing] extended the DMS method [dolbeault2015hypocoercivity] to analyze the ergodicity of Lindblad dynamics with broken detailed balance. In this work, we shall continue the investigation of the mixing property of Lindblad dynamics. We consider a general Lindbladian of the form:
$\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}\,,$
for a coupling parameter $\alpha>0$, where $\mathcal{L}^{H}=i[H,\cdot]$ is the coherent part and $\mathcal{L}^{D}$ is a Lindbladian satisfying $\sigma$-detailed balance. We assume that $\mathcal{L}_{\alpha}$ is primitive with $\sigma$ being a unique invariant state, which is of particular interest from the state preparation perspective. The related definitions and preliminaries for Lindblad dynamics are given in Section 2.1. We connect the spectral gap and the singular value gap of $\mathcal{L}_{\alpha}$ with its relaxation time in Section 2.2. In particular, we distinguish between the coercive and hypocoercive Lindblad dynamics by the primitivity of $\mathcal{L}^{D}$: ${\rm dim}\ker(\mathcal{L}^{D})=1$ (coercivity); ${\rm dim}\ker(\mathcal{L}^{D})>1$ (hypocoercivity); see Definition 2.4 and Remark 4.
While the convergence of coercive dynamics is well understood via Poincaré or modified log-Sobolev inequalities, our primary focus is on the hypocoercive case. In Section 3, we extend the variational framework [albritton2019variational, cao2023explicit, brigati2023construct] to explicitly quantify the exponential decay of the primitive hypocoercive Lindblad dynamics. The key step is to establish a quantum analog of space-time Poincaré inequality associated with $\mathcal{L}_{\alpha}$ for time-dependent operators (Theorem 3.2), relying on a technical abstract divergence lemma (Lemma 3.7). This space-time inequality corresponds to a pure exponential decay in the time-averaged $L^{2}$-distance (Theorem 3.1), revealing that for hypocoercive dynamics, one should consider a time interval rather than a single time slice to allow dissipation to propagate from the subspace to the whole space (Remark 5). We emphasize that the convergence rate estimate in Theorem 3.1 is explicit so that one can carefully choose the model parameters to optimize the rate; see Section 3.2. As applications of results in Section 3, some examples of Davies generators with broken detailed balance by a coherent term are provided in Section 5.
It is well known that the spectral gap provides the sharp $L^{2}$-convergence rate (see Section 2.2). To complement the hypocoercive rate estimate in Section 3, inspired by [hwang2005accelerating, franke2010behavior], we characterize the limiting behavior of the spectral gap $\lambda(\mathcal{L}_{\alpha})$ as $\alpha\to\infty$ (Theorem 4.1). We also found that under some mild assumption, adding a coherent term can strictly increase the spectral gap and hence accelerate the mixing of the Lindblad dynamics (Corollary 4.2). A very interesting but challenging question here is to construct the optimal Hamiltonian in the sense that the spectral gap is maximized, i.e., find the optimal value and the optimizer for $\sup\{\lambda(\mathcal{L}_{\alpha})\,;[H,\sigma]=0\}$ for a fixed $\alpha$.
Notation
[…]
2. Quantum Markov semigroup and its mixing
In Section 2.1 we introduce the detailed balanced quantum Markov semigroup (QMS) with a coherent term, as a special class of irreversible QMS, which would be the primary focus of this work. Then in Section 2.2, we discuss several concepts for analyzing the mixing property of QMS without detailed balance conditions.
2.1. Structures of quantum Markov semigroup
We first recall preliminaries for Markovian open quantum dynamics. A quantum Markov semigroup (QMS) $(\mathcal{P}_{t})_{t\geq 0}:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H})$ is a semigroup of completely positive and unital maps. Its generator $\mathcal{L}$, also called Lindbladian, is defined by
$\mathcal{L}(X):=\lim_{t\to 0}t^{-1}(\mathcal{P}_{t}(X)-X)\,,$
which has the following GKSL canonical form [Lindblad1976, GoriniKossakowskiSudarshan1976]:
(2.1) $\mathcal{L}(X)=i[H,X]+\sum_{j\in\mathcal{J}}\left(L_{j}^{\dagger}XL_{j}-\frac{1}{2}\left\{L_{j}^{\dagger}L_{j},X\right\}\right),$
with $i[H,X]$ and $\mathcal{L}(X)-i[H,X]$ being its coherent and dissipative parts, respectively. Here, the operator $H$ is the Hamiltonian while $\{L_{j}\}_{j\in\mathcal{J}}$ are called jump operators.
Let $\sigma\in\mathcal{D}_{+}(\mathcal{H})$ be a given full-rank quantum state. We define the modular operator:
$\Delta_{\sigma}(X)=\sigma X\sigma^{-1}\quad\text{for $X\in\mathcal{B}(\mathcal{H})$}\,,$
and the following family of inner products on $\mathcal{B}(\mathcal{H})$ with $s\in\mathbb{R}$: for any $X,Y\in\mathcal{B}(\mathcal{H})$,
(2.2) $\langle X,Y\rangle_{\sigma,s}:=\operatorname{tr}(\sigma^{s}X^{*}\sigma^{1-s}Y)=\langle X,\Delta_{\sigma}^{1-s}(Y)\sigma\rangle\,,$
where $\langle\cdot,\cdot\rangle_{\sigma,1}$ and $\langle\cdot,\cdot\rangle_{\sigma,1/2}$ are the GNS and KMS inner products, respectively. Then, a QMS $\mathcal{P}_{t}=e^{t\mathcal{L}}$ is said to satisfy the $\sigma$-GNS (resp., $\sigma$-KMS) detailed balance condition (DBC) if the Lindbladian $\mathcal{L}$ is self-adjoint with respect to the inner product $\langle\cdot,\cdot\rangle_{\sigma,1}$ (resp., $\langle\cdot,\cdot\rangle_{\sigma,1/2}$). In analogy with the classical case, when $\mathcal{L}$ is detailed balanced, we also say that the QMS $\mathcal{P}_{t}$ is reversible if there is no confusion from the context.
We recall the $\sigma$-weighted $p$-norm with $p\geq 1$ [olkiewicz1999hypercontractivity]: for $X\in\mathcal{B}(\mathcal{H})$,
$\lVert X\rVert_{p,\sigma}:=\operatorname{tr}\Big{(}|\Gamma_{\sigma}^{1/p}(X)|^{p}\Big{)}^{1/p}\,,$
where $|X|:=\sqrt{X^{*}X}$ is the modulus of $X\in\mathcal{B}(\mathcal{H})$. Noting $\lVert X\rVert_{2,\sigma}^{2}=\operatorname{tr}(|\Gamma_{\sigma}^{1/2}(X)|^{2})=\langle X,X\rangle_{\sigma,1/2}$, in what follows, we shall adopt the notation $\lVert\cdot\rVert_{\sigma,1/2}$ from the KMS inner product for the weighted $2$-norm, instead of $\lVert\cdot\rVert_{2,\sigma}$, for the sake of consistency. We also introduce the space of operators with expectation zero for $\sigma$:
(2.3) =$\mathfrak{H}=\left\{X\in\mathcal{B}(\mathcal{H})\,;\ \operatorname{tr}(\sigma X)=0\right\},$
and there holds $\mathcal{B}(\mathcal{H})=\operatorname{Ran}({\bf 1})\oplus\mathfrak{H}$. Then, the operator norm of a superoperator $\Phi$ on $\mathcal{B}(\mathcal{H})$ and $\mathfrak{H}$ are defined by
(2.4) $\lVert\Phi\rVert_{2\to 2}=\sup_{X\in\mathcal{B}(\mathcal{H})\backslash\{0\}}\frac{\lVert\Phi(X)\rVert_{\sigma,1/2}}{\lVert X\rVert_{\sigma,1/2}}\quad\text{and}\quad\lVert\Phi\rVert_{\mathfrak{H}\to\mathfrak{H}}=\sup_{X\in\mathfrak{H}\backslash\{0\}}\frac{\lVert\Phi(X)\rVert_{\sigma,1/2}}{\lVert X\rVert_{\sigma,1/2}}\,,$
respectively. To avoid confusion with the adjoint $\Phi^{\dagger}$ for Hilbert-Schmidt inner product, the adjoint of $\Phi$ for the KMS inner product $\langle\cdot,\cdot\rangle_{\sigma,1/2}$ is denoted by
$\Phi^{\star}=\Gamma_{\sigma}^{-1}\Phi^{\dagger}\Gamma_{\sigma}\,,$
where $\Gamma_{\sigma}$ is the weighting operator associated with $\sigma$:
$\Gamma_{\sigma}X:=\sigma^{\frac{1}{2}}X\sigma^{\frac{1}{2}}\,.$
Moreover, we introduce the Dirichlet form corresponding to a Lindbladian $\mathcal{L}$ by
$\mathcal{E}_{\mathcal{L}}(X,Y):=-\langle X,\mathcal{L}Y\rangle_{\sigma,1/2}\,,\quad\forall\,X,Y\in\mathcal{B}(\mathcal{H})\,.$
The next proposition shows that the KMS detailed balanced Lindbladians can be written as squares of derivations in some sense [vernooij2023derivations]*Theorem 2.5.
Proposition 2.1.
Let $\mathcal{L}$ be a Lindbladian satisfying $\sigma$-KMS DBC. Then there exist $\{V_{j}\}_{j=1}^{\mathcal{J}_{D}}\subset\mathcal{B}(\mathcal{H})$ such that the associated Dirichlet form can be written as
$\mathcal{E}_{\mathcal{L}}(X,Y)=\sum_{j=1}^{\mathcal{J}_{D}}\langle[V_{j},X],[V_{j},Y]\rangle_{\sigma,1/2}\,,$
where $[A,B]=AB-BA$ denotes the commutator.
In this work, we focus on the following family of irreversible QMS: for $\alpha>0$,
(2.5) $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}\,,$
where $\alpha>0$ is the coupling parameter, representing the relative strength between $\mathcal{L}^{H}$ and $\mathcal{L}^{D}$. Here, the dissipative part $\mathcal{L}^{D}$ is a Lindbladian satisfying $\sigma$-KMS DBC for some $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, and the coherent term $\mathcal{L}^{H}(\cdot):=i[H,\cdot]$ is defined by a Hamiltonian $H$ satisfying $[H,\sigma]=0$. It follows that $\sigma$ is a fixed point of the Lindblad dynamic generated by $\mathcal{L}^{\dagger}_{\alpha}$, i.e.,
$\mathcal{P}_{t}^{\dagger}(\sigma)=e^{t\mathcal{L}_{\alpha}^{\dagger}}(\sigma)=\sigma\,,\quad\text{equivalently},\quad\mathcal{L}_{\alpha}^{\dagger}(\sigma)=0\,.$
We shall limit our discussion to the case where the QMS $\mathcal{P}_{t}$ is primitive, that is, $\sigma$ is the unique invariant state. In this case, there holds [frigerio1982long]
(2.6) $\lim_{t\to\infty}\mathcal{P}_{t}(X)=\operatorname{tr}(\sigma X){\bf 1}\,,\quad\forall\,X\in\mathcal{B}(\mathcal{H})\,.$
It is worth noting, by $[H,\sigma]=0$,
$\langle Y,[H,X]\rangle_{\sigma,1/2}=\langle Y,H\sigma^{1/2}X\sigma^{1/2}-\sigma^{1/2}X\sigma^{1/2}H\rangle=\langle[H,Y],X\rangle_{\sigma,1/2}\,,$
which implies that $\mathcal{L}^{H}=i[H,\cdot]$ is anti-Hermitian for the KMS inner product:
$\mathcal{L}^{H}=-(\mathcal{L}^{H})^{\star}\,,$
resulting in $\mathcal{L}_{\alpha}$ being irreversible. However, we will see in Remark 2 that the reversibility (detailed balance) of $\mathcal{L}_{\alpha}$ can be recovered in a generalized sense. The following lemma gives useful characterizations of $\ker(\mathcal{L})$ and the primitivity of the Lindbladian.
Lemma 2.2 ([wolf5quantum]*Theorem 7.2).
Suppose that the QMS $\mathcal{P}^{\dagger}_{t}=e^{t\mathcal{L}^{\dagger}}$ admits a full-rank invariant state $\sigma$, namely, $\mathcal{L}^{\dagger}(\sigma)=\sigma$, where $\mathcal{L}$ is of the form (2.1). Then,
$\{H,L_{j},L_{j}^{\dagger}\}^{\prime}=\ker(\mathcal{L})\,,$
and hence the primitivity is also equivalent to
(2.7) $\{H,L_{j},L_{j}^{\dagger}\}^{\prime}=\{z{\bf 1}\,;\ z\in\mathbb{C}\}\,,$
where the commutant $\{H,L_{j},L_{j}^{\dagger}\}^{\prime}$ is defined by the operators that commute with $L_{j}$, $L_{j}^{\dagger}$ and $H$.
Let $\{V_{j}\}_{j\in\mathcal{J}_{D}}$ be the operators associated with $\mathcal{L}^{D}$ given by Proposition 2.1. We define the noncommutative partial derivative by
$\partial_{j}X=[V_{j},X]\,,\quad\text{with}\quad\partial_{j}^{\dagger}X=[V_{j}^{*},X]\,,$
and the associated gradient $\nabla:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H})^{\mathcal{J}_{D}}$ by
$\nabla X=(\partial_{1}X,\cdots,\partial_{J}X)\,,\quad\text{for}\ X\in\mathcal{B}(\mathcal{H})\,.$
This allows us to write $\mathcal{E}_{\mathcal{L}^{D}}(X,Y)=\sum_{j}\langle\partial_{j}X,\partial_{j}Y\rangle_{\sigma,1/2}$ and further obtain
(2.8) $\mathcal{L}^{D}(X)=-\sum_{j\in\mathcal{J}_{D}}\partial_{j}^{\star}\partial_{j}X=-\nabla^{\star}\nabla X\,.$
Here, $\partial_{j}^{\star}$ is the adjoint of $\partial_{j}$ with respect to $\langle\cdot,\cdot\rangle_{\sigma,1/2}$:
(2.9) $\partial_{j}^{\star}X=\Gamma_{\sigma}^{-1}\partial_{j}^{\dagger}\Gamma_{\sigma}X=\Gamma_{\sigma}^{-1}[V_{j}^{*},\Gamma_{\sigma}X]\,.$
Then, one can readily reformulate the Lindbladian $\mathcal{L}_{\alpha}$ in Eq. 2.5 into the hypocoercive form:
(2.10) $\mathcal{L}_{\alpha}(X)=i\alpha[H,X]-\nabla^{\star}\nabla X\,.$
which can be regarded as a quantum analog of underdamped Langevin dynamics; see Remark 1. The following lemma is a direct consequence of (2.8) and (2.10).
Lemma 2.3.
For Lindbladian $\mathcal{L}^{D}$ in (2.8), there holds
$\ker(\mathcal{L}^{D})=\ker(\nabla)=\{X\,;\ [X,V_{j}]=0\ \text{forall}\ V_{j}\,,\,j\in\mathcal{J}_{D}\}\,.$
For $\mathcal{L}_{\alpha}$ in (2.10), there holds
$\ker(\mathcal{L}_{\alpha})=\ker(\mathcal{L}^{D})\bigcap\ker(\mathcal{L}^{H})\,.$
Remark 1 (Quantum analog of kinetic Fokker-Planck equation).
We recall some basics of the classical hypocoercive dynamics to better appreciate the connection of Lindblad dynamics in the form (2.10) with kinetic Fokker-Planck equations. Let $L$ be a generator for a Markov process with $\mu$ being the unique invariant measure. The Markov evolution $e^{tL}$ is in the hypocoercive form if
(2.11) $L=B-A^{*}A\,,$
for some operators $A,B$, where $B=-B^{*}$ is anti-symmetric. Here $A^{*}$ and $B^{*}$ denote the adjoints with respect to $\langle\cdot,\cdot\rangle_{2,\mu}$. We refer the readers to [villani2009hypocoercivity]*Section 3 for the rigorous definition of a hypocoercive operator. A well-studied hypocoercive model is the kinetic Fokker-Planck equation: given the potential energy $U(x)$ and the friction coefficient $\gamma>0$,
$\partial_{t}\rho(t,x,v)=-v\cdot\nabla_{x}\rho+\nabla_{x}U\cdot\nabla_{v}\rho+\gamma(\Delta_{v}+\nabla_{v}\cdot(v\rho))\,,$
where $\rho(t,\cdot)$ denotes a distribution on $(x,v)\in\mathbb{R}^{2d}$. Under mild assumptions, it has a unique invariant probability measure:
$\mathrm{d}\rho_{\infty}(x,v)=\,\mathrm{d}\mu(x)\,\mathrm{d}\kappa(v)\,,$
where $\mathrm{d}\mu=e^{-U(x)}\,\mathrm{d}x/Z_{U}$ and $\mathrm{d}\kappa=e^{-|v|^{2}/2}\,\mathrm{d}v/(2\pi)^{d/2}$. Its corresponding backward Kolmogorov equation has the generator $\mathcal{L}_{\rm kFP}=\mathcal{L}_{\rm ham}+\gamma\mathcal{L}_{\rm FD}$ with
$\mathcal{L}_{\rm ham}:=\nabla_{v}^{*}\nabla_{x}-\nabla_{x}^{*}\nabla_{v}\,,\quad\mathcal{L}_{\rm FD}:=-\nabla_{v}^{*}\nabla_{v}\,.$
Here $\nabla_{x}^{*}$ and $\nabla_{v}^{*}$ are the adjoints of $\nabla_{x}$ and $\nabla_{v}$ for the inner product induced by $\rho_{\infty}$.
By above discussions, one can readily see that both $\mathcal{L}_{\alpha}$ and $\mathcal{L}_{\rm kFP}$ can fit into the form (2.11), and $\mathcal{L}_{\alpha}$ can be naturally viewed as a quantum analog of $\mathcal{L}_{\rm kFP}$. We would also like to mention the recent work [galkowski2024classical] on the correspondence between the Lindblad and Fokker-Planck evolutions, which shows that if $H$ and $L_{j}$ in (2.1) are given by the semiclassical quantizations of classical observables, then the leading part of the semiclassical expansion of $\mathcal{P}_{t}$ gives the Fokker-Planck dynamics.
Remark 2 (Generalized detailed balance and time reversal).
Let $\mu$ be a probability on $\mathbb{R}^{d}$ and $\mathcal{F}$ be an involution (i.e., a bounded operator with $\mathcal{F}^{2}=1$) that leaves $\mu$ invariant: $\mu\circ\mathcal{F}=\mu$. A Markov semigroup $e^{t\mathcal{L}}$ is generalized reversible with respect to $\mu$ up to $\mathcal{F}$ if and only if
(2.12) $\mathcal{L}^{*}=\mathcal{L}(f\circ\mathcal{F})\circ\mathcal{F}\,,$
where $\mathcal{L}^{*}$ is the adjoint of $\mathcal{L}$ for $L^{2}_{\mu}$. When $\mathcal{F}$ is the identity, the above definition reduces to the usual reversibility. One can show that the generator $\mathcal{L}_{\rm kFP}$ for the kinetic Fokker-Planck equation in Remark 1 is irreversible in the usual sense but satisfies the generalized reversibility with respect to $\rho_{\infty}$ up to $\mathcal{F}(x,v)=(x,-v)$ [stoltz2010free].
The quantum DBC (reversibility) can be generalized in various ways, which is typically formulated by a certain relationship between the Lindbladian $\mathcal{L}$ and its adjoint with respect to some inner product. The discussion below focuses on the KMS inner product, following [fagnola2010generators]; a similar discussion with the GNS inner product can be found in [fagnola2008detailed] and references therein. A Lindbladian satisfies the standard DBC if there exists a self-adjoint operator $K$ such that [derezinski2006fermi]
(2.13) $\mathcal{L}-\mathcal{L}^{\star}=2i[K,\cdot]\,,$
where $\mathcal{L}^{\star}$ is the adjoint of $\mathcal{L}$ for the KMS inner product. It is clear that $\mathcal{L}_{\alpha}$ introduced in (2.10) is detailed balanced in the sense of (2.13) with $K=H$. Another generalization of quantum DBC involves a time reversal $T$ and a similar formulation as the classical one (2.12). Here a time reversal $T$ means an antiunitary operator on $\mathcal{H}$ (i.e., $\langle Tx,Ty\rangle=\langle y,x\rangle$ for $x,y\in\mathcal{H}$) such that $T^{-1}=T$. We say a Lindbladian satisfies the standard DBC with time reversal $T$ if
(2.14) $\mathcal{L}^{\star}(X)=T\mathcal{L}(TXT^{-1})T^{-1}\,.$
Unfortunately, in general, the definitions (2.13) and (2.14) are not comparable. We refer the interested readers to [guo2024designing] for further connections between the classical and quantum DBC.
2.2. Spectral gap and relaxation time
Let $\mathcal{L}$ be a primitive Lindbladian that admits a full-rank state $\sigma\in\mathcal{D}_{+}(\mathcal{H})$ as an invariant state. We will discuss various measures for quantifying the mixing properties of the Lindblad dynamics $\mathcal{P}_{t}=e^{t\mathcal{L}}$.
We first introduce the spectral gap of $\mathcal{L}$:
(2.15) $\lambda(\mathcal{L})=\inf\left\{\mathfrak{Re}(\lambda)\,;\ \lambda\in{\rm Spec}(-\mathcal{L}|_{\mathfrak{H}})\right\},$
where the subspace $\mathfrak{H}$ is defined in (2.3) and $\mathcal{L}|_{\mathfrak{H}}$ is the restriction of the Lindbladian $\mathcal{L}$ on $\mathfrak{H}$. For $\varepsilon\in(0,1)$, the non-asymptotic $\varepsilon$-relaxation time of $\mathcal{P}_{t}$ in the $L^{2}$-sense is given by
$t_{\rm rel}(\varepsilon):=\inf\left\{t\geq 0\,;\ \lVert\mathcal{P}_{t}X\rVert_{\sigma,1/2}\leq\varepsilon\lVert X\rVert_{\sigma,1/2}\ \,\text{for all}\ \,X\in\mathfrak{H}\right\},$
which, as a function $\varepsilon\to t_{\rm rel}(\varepsilon)$, is the inverse function of $t\to\lVert\mathcal{P}_{t}\rVert_{\mathfrak{H}\to\mathfrak{H}}$. In the case of $\mathcal{L}$ satisfying $\sigma$-KMS DBC, it is easy to see
(2.16) $\lVert\mathcal{P}_{t}\rVert_{\mathfrak{H}\to\mathfrak{H}}=e^{-\lambda(\mathcal{L})t}\,.$
This implies that the relaxation time is proportional to the inverse of the spectral gap:
$t_{\rm rel}(\varepsilon)=\log(\varepsilon^{-1})/\lambda(\mathcal{L})\propto\lambda(\mathcal{L})^{-1}\,.$
When the detailed balance of $\mathcal{L}$ is broken, the pure exponential decay in (2.16) does not hold and one can only expect
(2.17) $\lVert\mathcal{P}_{t}\rVert_{\mathfrak{H}\to\mathfrak{H}}\leq Ce^{-\nu t}\,,$
for some $C\in[1,\infty)$ and $0<\nu\leq\lambda(\mathcal{L})$. We define the sharp exponential decay rate by
[…]
By the semigroup theory of linear evolution equations [engel2000one]*Chapter IV, we have
$\lambda(\mathcal{L})=\nu_{0}=-\lim_{t\to\infty}\frac{1}{t}\log\lVert\mathcal{P}_{t}\rVert_{\mathfrak{H}\to\mathfrak{H}}\,,$
which can be equivalently formulated as
$\lambda(\mathcal{L})^{-1}=\lim_{\varepsilon\to 0}\frac{t_{\rm rel}(\varepsilon)}{\log(\varepsilon^{-1})}\,.$
It means that in the asymptotic regime $\varepsilon\to 0$, we still have $t_{\rm rel}(\varepsilon)\propto\lambda(\mathcal{L})^{-1}$, while for finite $\varepsilon>0$, the inverse spectral gap only gives a lower bound of the relaxation time:
(2.19) $t_{\rm rel}(\varepsilon)\geq\log(\varepsilon^{-1})/\lambda(\mathcal{L})\,,$
due to the existence of the multiplicative constant $C\geq 1$ in (2.17).
Remark 3.
The supremum $\nu_{0}$ in (2.18) may not be attained in the general case. To be precise, given a primitive Lindblad dynamics $\mathcal{P}_{t}$, one can only have [^1] This happens when $\mathcal{L}$ has an eigenvalue $\lambda\neq 0$ such that $\lambda(\mathcal{L})=\mathfrak{Re}(-\lambda)$ and its geometric multiplicity is strictly less than its algebraic multiplicity., for any $\epsilon>0$,
$\lVert\mathcal{P}_{t}\rVert_{\mathfrak{H}\to\mathfrak{H}}\leq C_{\epsilon}e^{-(\lambda(\mathcal{L})-\epsilon)t}\quad\text{for some $C_{\epsilon}\geq 1$}\,,$
see [arnold2014sharp]*Theorem 4.9. The sharp decay rate $\nu_{0}=\lambda(\mathcal{L})$ can be recovered by allowing a time-polynomial prefactor $C(t)$: $\lVert\mathcal{P}_{t}\rVert_{\mathfrak{H}\to\mathfrak{H}}\leq C(t)e^{-\lambda(\mathcal{L})t}$ [arnold2020sharp]*Theorem 2.8.
Following the technology in kinetic theory [villani2009hypocoercivity], we introduce the following definition.
Definition 2.4.
A primitive Lindblad dynamics $\mathcal{P}_{t}=e^{t\mathcal{L}}$ is hypocoercive if the estimate (2.17) holds for some $C>1$ and $\nu>0$, and $\mathcal{P}_{t}$ is called coercive if (2.17) holds with $C=1$.
Remark 4 (Coercivity and symmetrized Lindbladian).
From Eqs. 2.16 and 2.17 with $C=1$, we readily see that the primitive KMS-detailed balanced Lindbladian $\mathcal{L}$ is coercive. However, the converse is generally not true. In fact, by taking the derivative of $\lVert\mathcal{P}_{t}X\rVert_{\sigma,1/2}\leq e^{-\nu t}\lVert X\rVert_{\sigma,1/2}$ for $X\in\mathfrak{H}$ and Grönwall’s inequality, we see that the coercivity of $\mathcal{L}$ is equivalent to the Poincaré inequality of the additively symmetrized Lindbladian: for $X\in\mathfrak{H}$,
$\mathcal{E}_{\frac{\mathcal{L}+\mathcal{L}^{\star}}{2}}(X,X)\geq\nu\lVert X\rVert_{\sigma,1/2}^{2}\,.$
Moreover, let $X$ be the eigenvector of $\mathcal{L}$ associated with the spectral gap $\lambda(\mathcal{L})$, that is, $\mathcal{L}X=\lambda X$ with $\mathfrak{Re}(\lambda)=-\lambda(\mathcal{L})$. We have $-\langle X,\frac{\mathcal{L}+\mathcal{L}^{\star}}{2}X\rangle_{\sigma,1/2}=\lambda(\mathcal{L})\lVert X\rVert_{\sigma,1/2}^{2}$, which implies
(2.20) $\lambda(\mathcal{L})\geq\inf_{X\in\mathfrak{H}\backslash\{0\}}\frac{\mathcal{E}_{\frac{\mathcal{L}+\mathcal{L}^{\star}}{2}}(X,X)}{\lVert X\rVert_{\sigma,1/2}^{2}}\,,$
meaning that the convergence rate from the coercivity of the symmetrized Lindbladian is generally not sharp. For the primitive model (2.5), there holds $\mathcal{L}_{\alpha}+\mathcal{L}_{\alpha}^{\star}=2\mathcal{L}^{D}$. In this case, the coercivity of $\mathcal{L}_{\alpha}$ is equivalent to that $\mathcal{L}^{D}$ is injective on $\mathfrak{H}$, i.e., $\dim\ker(\mathcal{L}^{D})=1$. In other words, $\mathcal{L}_{\alpha}$ is hypocoercive if and only if $\dim\ker(\mathcal{L}^{D})>1$ (see A).
To establish an upper bound on the relaxation time $t_{\rm rel}(\varepsilon)$, we reformulate (2.17) as the following $T$-delayed version: for some $\nu>0$ and $T\geq 0$,
(2.21) $\lVert\mathcal{P}_{t}X\rVert_{\sigma,1/2}\leq e^{-\nu(t-T)}\lVert X\rVert_{\sigma,1/2}\,,\quad\forall\,X\in\mathfrak{H}\,,$
which can be implied by the $T$-average exponential decay with rate $\nu$ (see the proof of Theorem 3.1):
(2.22) $\frac{1}{T}\int_{t}^{t+T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s\leq e^{-\nu t}\frac{1}{T}\int_{0}^{T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s\,,\quad\forall\,X\in\mathfrak{H}\,.$
With the above notions, it is standard to define, for $\varepsilon\in(0,1)$,
[…]
which is of the same order as $t_{\rm rel}(\varepsilon)$:
(2.23) $t_{\rm rel}(\varepsilon)\leq t_{0}(\varepsilon)\leq 2t_{\rm rel}(\varepsilon)\,.$
Indeed, the left-hand side inequality of (2.23) is direct. Recalling (2.19), the right-hand side one is by taking $t=2t_{\rm rel}(\varepsilon)$, $T=t_{\rm rel}(\varepsilon)$ and $\nu=\log(\varepsilon^{-1})/t_{\rm rel}(\varepsilon)$ in (2.21):
$\lVert\mathcal{P}_{t}X\rVert_{\sigma,1/2}\leq e^{-\nu t_{\rm rel}(\varepsilon)}\lVert X\rVert_{\sigma,1/2}=\varepsilon\lVert X\rVert_{\sigma,1/2}\,.$
Therefore, to bound $t_{\rm rel}(\varepsilon)$ from above, it suffices to find a lower bound of $\nu$ for a fixed $T$ such that the averaged decay (2.22) holds, which is the main purpose of Section 3.
The singular value estimate, corresponding to the multiplicative symmetrization of the generator, has been widely used for analyzing the mixing property of a classical irreversible Markov process [fill1991eigenvalue, chatterjee2023spectral]. Temme et al. [temme2010chi] characterized the convergence of the discrete-time quantum Markov process in the $\chi^{2}$-divergence by the singular value gap of the quantum channel. Here, we briefly discuss the connections between the singular value gap of $\mathcal{L}$ and the relaxation time, which extends [eberle2024non]*Lemma 10. We define
$s(\mathcal{L}):=\lambda\Big{(}\sqrt{\mathcal{L}^{\star}\mathcal{L}}\Big{)}=\inf_{X\in\mathfrak{H}\backslash\{0\}}\frac{\lVert\mathcal{L}X\rVert_{\sigma,1/2}}{\lVert X\rVert_{\sigma,1/2}}\,,$
and we have $|\lambda|\geq s(\mathcal{L})$ for any $\lambda\in{\rm Spec}(\mathcal{L}|_{\mathfrak{H}})$. It is also clear that
$\frac{1}{s(\mathcal{L})}=s\left(\mathcal{L}^{-1}\right)=\inf_{X\in\mathfrak{H}\backslash\{0\}}\frac{\lVert\mathcal{L}^{-1}X\rVert_{\sigma,1/2}}{\lVert X\rVert_{\sigma,1/2}}\,.$
Lemma 2.5.
Suppose that $\mathcal{P}_{t}=e^{t\mathcal{L}}$ is a primitive Lindblad dynamics satisfying the $T$-average exponential decay with rate $\nu$. Then it holds that $\frac{1}{\nu}+T\geq\frac{1}{s(\mathcal{L})}$, and hence $t_{\rm rel}(e^{-1})\geq\frac{1}{2s(\mathcal{L})}$.
Proof.
Note that the $T$-average exponential decay yields the $2T$-average one with the same rate:
$\frac{1}{2T}\int_{t}^{t+2T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s$ $\leq\frac{1}{2T}\int_{t}^{t+T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s+\frac{1}{2T}\int_{t+T}^{t+2T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s$
$\leq e^{-\nu t}\frac{1}{2T}\int_{0}^{2T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s\,,\quad\forall\,X\in\mathfrak{H}\,.$
Then, by the $2T$-average exponential decay, we have, for $X\in\mathfrak{H}$,
$\int_{0}^{\infty}\lVert\mathcal{P}_{t}X\rVert_{\sigma,1/2}\,\mathrm{d}t$ $=\int_{0}^{\infty}\frac{1}{2T}\int_{t}^{t+2T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}\,\mathrm{d}s\,\mathrm{d}t+\frac{1}{2T}\int_{0}^{2T}(2T-t)\lVert\mathcal{P}_{t}X\rVert_{\sigma,1/2}\,\mathrm{d}t$
$\leq\left(\int_{0}^{\infty}e^{-\nu t}\,\mathrm{d}t+\frac{1}{2T}\int_{0}^{2T}(2T-t)\,\mathrm{d}t\right)\lVert X\rVert_{\sigma,1/2}$
$\leq\left(\frac{1}{\nu}+T\right)\lVert X\rVert_{\sigma,1/2}\,.$
It follows that the operator $\int_{0}^{\infty}\mathcal{P}_{t}X\,\mathrm{d}t$ is well-defined, which is the inverse of the operator $-\mathcal{L}:\mathfrak{H}\to\mathfrak{H}$. Moreover, the following estimate holds:
$\lVert\mathcal{L}^{-1}(X)\rVert_{\sigma,1/2}\leq\left(\frac{1}{\nu}+T\right)\lVert X\rVert_{\sigma,1/2}\,,$
which implies $1/s(\mathcal{L})\leq 1/\nu+T$. The proof is completed by (2.23). ∎
3. Convergence rate estimate via hypocoercivity
In this section, we provide an explicit and constructive decay rate estimate of the exponential convergence of the hypocoercive primitive QMS (2.10) in the $L^{2}$-distance and the time-averaged one. This analysis is motivated by the variational framework based on space-time Poincaé inequality, recently developed in [albritton2019variational, cao2023explicit, brigati2023construct] for kinetic Fokker-Planck equations.
3.1. Main result
For ease of exposition, we set the coupling parameter $\alpha=1$ in Eq. 2.10 (as we can always absorb $\alpha>0$ in $\mathcal{L}^{H}$) and consider the primitive Lindbladian:
(3.1) $\mathcal{L}=\mathcal{L}^{H}+\mathcal{L}^{D}\,,$
where $\mathcal{L}^{H}=i[H,\cdot]$ and $\mathcal{L}^{D}$ is a $\sigma$-KMS detailed balanced Lindbladian. Our main result is the decay estimate in Theorem 3.1, which is based on a novel space-time Poincaré inequality for time-dependent operators (see Theorem 3.2). The scaling of the convergence rate in various parameters, the optimal selection of $\alpha$, and the comparison with the recent work [fang2024mixing] will be discussed in Section 3.2.
We recall the space $\mathfrak{H}$ of operators with average zero (2.3). Let $\Pi_{0}$ be the orthogonal projection (i.e., the conditional expectation [bardet2017estimating]) to $\ker({\mathcal{L}^{D}|_{\mathfrak{H}}})$ with respect to $\langle\cdot,\cdot\rangle_{\sigma,1/2}$. For simplicity of notation, we will often abuse the notation and suppress the restriction to $\mathfrak{H}$ in the sequel. We denote by $\Pi_{+}=1-\Pi_{0}$ the orthogonal complement of $\Pi_{0}$, and then $\mathfrak{H}$ can be decomposed as
(3.2) $\mathfrak{H}=\mathfrak{H}_{0}\oplus\mathfrak{H}_{+}\,,$
where
$\mathfrak{H}_{0}:=\ker(\mathcal{L}^{D}|_{\mathfrak{H}})\quad\text{and}\quad\mathfrak{H}_{+}:=\operatorname{Ran}(\Pi_{+})=\ker(\mathcal{L}^{D}|_{\mathfrak{H}})^{\perp}\,.$
Thanks to the finite dimensionality of $\mathcal{H}$ and the dissipativity of $\mathcal{L}^{D}$ with detailed balance condition, the Poincaré inequality for $\mathcal{L}^{D}$ holds on $\mathfrak{H}_{+}$: there exists a positive constant $\lambda_{D}>0$ such that
(3.3) $\mathcal{E}_{\mathcal{L}^{D}}(X,X)=-\langle X,\mathcal{L}^{D}X\rangle_{\sigma,1/2}\geq\lambda_{D}\lVert X\rVert_{\sigma,1/2}^{2}\,,\quad\forall\,X\in\mathfrak{H}_{+}\,,$
where $\lambda_{D}$ characterizes the spectral gap of $\mathcal{L}^{D}$. It follows that
(3.4) $\mathfrak{Re}\mathcal{E}_{\mathcal{L}}(X,X)=-\mathfrak{Re}\langle X,(\mathcal{L}^{H}+\mathcal{L}^{D})X\rangle_{\sigma,1/2}\geq\lambda_{D}\lVert X\rVert_{\sigma,1/2}^{2}\,,\quad\forall\,X\in\mathfrak{H}_{+}\,,$
since $\mathcal{L}^{H}$ is anti-Hermitian for $\langle\cdot,\cdot\rangle_{\sigma,1/2}$.
Remark 5.
The subspaces $\mathfrak{H}_{0}$ and $\mathfrak{H}_{+}$ correspond to the non-decaying and decaying modes of the Lindblad dynamics $e^{t\mathcal{L}^{D}}$, respectively. For the entire dynamics $e^{t\mathcal{L}}$ which is primitive, the Hamiltonian flow $e^{t\mathcal{L}^{H}}$ facilitates the propagation of the dissipation on $\mathfrak{H}_{+}$ to $\mathfrak{H}_{0}$ (and hence the whole space $\mathfrak{H}$), ensuring the convergence of $e^{t\mathcal{L}}$ to the equilibrium. However, the standard energy estimate, based on the Poincaré inequality (3.4), only implies the naive $L^{2}$-contraction, rather than the exponential decay, unless $\mathfrak{H}_{+}=\mathfrak{H}$ (see Remark 4). It means that new proof techniques are needed. In particular, we shall consider the time-averaged Lindblad dynamics (2.22), instead of a single time slice (i.e., $T\to 0$ for $\frac{1}{T}\int_{t}^{t+T}\lVert\mathcal{P}_{s}X\rVert_{\sigma,1/2}$); see also Remark 7.
To estimate the exponential decay rate of Lindblad dynamics (3.1), we introduce the following assumption, inspired by the hypocoercivity theory of classical kinetic equations. See Remark 1 and Remark 6 for detailed discussions.
Assumption A.
The subspace $\ker({\mathcal{L}^{D}|_{\mathfrak{H}}})$ is non-trivial, i.e., $\dim\ker({\mathcal{L}^{D}|_{\mathfrak{H}}})\geq 1$. In addition, the coherent term satisfies
(3.5) $\Pi_{0}\mathcal{L}^{H}\Pi_{0}=0\,.$
With the help of the decomposition (3.2) and A, the generator $\mathcal{L}$, restricted on the space $\mathfrak{H}$, can be reformulated into the following block form on $\mathfrak{H}_{0}\oplus\mathfrak{H}_{+}$:
(3.6) $\mathcal{L}|_{\mathfrak{H}}=\left[\begin{matrix}0&\mathcal{L}^{H}_{0+}\\ \mathcal{L}^{H}_{+0}&\mathcal{L}_{++}\end{matrix}\right],$
where
+$\mathcal{L}^{H}_{+0}:=\Pi_{+}\mathcal{L}^{H}\Pi_{0}=\mathcal{L}^{H}\Pi_{0}\quad\text{and}\quad\mathcal{L}^{H}_{0+}:=\Pi_{0}\mathcal{L}^{H}\Pi_{+}=-(\mathcal{L}^{H}_{+0})^{\star}\,,$
and by the inequality (3.4), $\mathcal{L}_{++}:=\Pi_{+}\mathcal{L}\Pi_{+}$ is invertible on $\mathfrak{H}_{+}$ with $\lVert\mathcal{L}_{++}^{-1}\rVert_{2\to 2}\leq\lambda_{D}^{-1}$.
Recalling that the QMS $\mathcal{P}_{t}$ is assumed to be primitive, it readily follows that $\mathcal{L}$ is invertible on $\mathfrak{H}$, and by the block form (3.6) under A, there necessarily holds
$\ker(\mathcal{L}^{H}_{+0})=\ker(\Pi_{0})\,,$
equivalently, +$\mathcal{L}^{H}_{+0}$ is injective on $\mathfrak{H}_{0}$. It means that (+$(\mathcal{L}^{H}_{+0})^{\star}\mathcal{L}^{H}_{+0}X=0$ for a $X\in\mathfrak{H}_{0}$ if and only if $X=0$. Thus, there exists $s_{H}>0$ such that for any $X\in\mathfrak{H}_{0}$,
(3.7) $\lVert\mathcal{L}_{+0}^{H}X\rVert_{\sigma,1/2}\geq s_{H}\lVert X\rVert_{\sigma,1/2}\,.$
Here, $s_{H}$ characterizes the singular value gap of $\mathcal{L}_{+0}^{H}$. Without loss of generality, we let $\lambda_{D}$ and $s_{H}$ be the largest (optimal) constants such that Eqs. 3.3 and 3.7 hold:
(3.8) $\lambda_{D}=\inf_{X\in\mathfrak{H}_{+}\backslash\{0\}}\frac{\mathcal{E}_{\mathcal{L}^{D}}(X,X)}{\lVert X\rVert_{\sigma,1/2}^{2}}\,,\quad s_{H}=\inf_{X\in\mathfrak{H}_{0}\backslash\{0\}}\frac{\lVert\mathcal{L}_{+0}^{H}X\rVert_{\sigma,1/2}}{\lVert X\rVert_{\sigma,1/2}}\,.$
Remark 6.
As has been discussed in Remark 4, the condition $\dim\ker({\mathcal{L}^{D}|_{\mathfrak{H}}})\geq 1$ in A is equivalent to the hypocoercivity of $\mathcal{L}$. Indeed, if $\ker({\mathcal{L}^{D}|_{\mathfrak{H}}})=0$, then $\Pi_{0}$ and $s_{H}$ have to be zero as well. In this case (i.e., $\dim\ker(\mathcal{L}^{D})=1$), the Lindbladian is coercive and the analysis in this section based on the space-time Poincaré inequality does not apply.
The condition (3.5) is motivated by [dolbeault2015hypocoercivity]*Assumption (H3) for linear kinetic equations; see also [bernard2022hypocoercivity]*Assumption 2.1. For classical dynamics, inequalities (3.3) and (3.7) are usually referred to as the microscopic and macroscopic coercivities, respectively, which are essentially the Poincaré inequalities of measures $\kappa$ and $\mu$. To be specific, let us again consider the kinetic Fokker-Planck equation recalled in Remark 1. For this model, we have $\Pi_{0}\varphi(x)=\int_{\mathbb{R}^{d}}\varphi(x,v)\,\mathrm{d}\kappa(v)$ and $\mathcal{L}^{D}=\gamma\mathcal{L}_{\rm FD}$, the inequality (3.3) can be formulated as: for some $\lambda_{v}>0$,
$-\gamma\lVert\nabla_{v}f\rVert_{2,\rho_{\infty}}^{2}\geq\lambda_{v}\lVert(1-\Pi_{0})f\rVert_{2,\rho_{\infty}}^{2}\,,$
which can be implied by the Poincaré inequality of the measure $\kappa(v)$ in the momentum variable $v$. In addition, a direct computation as in [bernard2022hypocoercivity]*Theorem 3.3 shows, by replacing $\mathcal{L}^{H}$ with $\mathcal{L}_{\rm ham}$,
(+$(\mathcal{L}^{H}_{+0})^{\star}\mathcal{L}^{H}_{+0}=\Pi_{0}\mathcal{L}_{\rm ham}^{*}\mathcal{L}_{\rm ham}\Pi_{0}=\sum_{i,j=1}^{d}M_{ij}\Pi_{0}\partial_{x_{i}}^{*}\partial_{x_{j}}\Pi_{0}\,,$
where $M$ is the moment matrix of $\kappa$: $M_{ij}=\int_{\mathbb{R}^{d}}v_{i}v_{j}\,\mathrm{d}\kappa(v)$. Then, by Poincaré inequality of $\mu(x)$ in the position variable $x$, we have, for some $\lambda_{x}>0$,
$\lVert\mathcal{L}_{\rm ham}\Pi_{0}f\rVert^{2}_{2,\rho_{\infty}}\geq\lambda_{\min}(M)\lVert\nabla_{x}\Pi_{0}f\rVert^{2}_{2,\mu}\geq\lambda_{\min}(M)\lambda_{x}\lVert\Pi_{0}f\rVert^{2}_{2,\mu}\,,$
that is, Eq. 3.7 holds. Here $\lambda_{\min}(M)$ denotes the minimal eigenvalue of $M$.
Before we state the main results, we introduce some additional notations that will be used in what follows. We consider the time interval $[0,T]$ with the normalized Lebesgue measure:
(3.9) $\mathrm{d}\lambda(t):=\frac{1}{T}\chi_{[0,T]}(t)\,\mathrm{d}t\,.$
We denote by $H_{\lambda}^{k}([0,T])$ with $k\geq 0$ the standard Sobolev space of functions $f(t)$ on $[0,T]$, and by $\langle f,g\rangle_{2,\lambda}=\int_{0}^{T}f^{*}g\,\mathrm{d}\lambda$ the inner product with respect to $\lambda$.
Given a full-rank quantum state $\sigma$, for a time-dependent operator $(X_{t})_{t\in[0,T]}\subset\mathcal{B}(\mathcal{H})$, we define its state average, time average, and the state-time average by,
$\langle X_{t}\rangle_{\sigma}:=\operatorname{tr}(\sigma X_{t})\,,\quad\langle X_{t}\rangle_{\lambda}:=\frac{1}{T}\int_{0}^{T}X_{t}\,\mathrm{d}t\,,\quad\langle X_{t}\rangle_{\lambda\otimes\sigma}:=\langle\langle X_{t}\rangle_{\sigma}\rangle_{\lambda}=\frac{1}{T}\int_{0}^{T}\operatorname{tr}(\sigma X_{t})\,\mathrm{d}t\,,$
respectively. Here $\lambda\otimes\sigma$ can be viewed as an augmented quantum state on the time-operator product space $[0,T]\times\mathcal{B}(\mathcal{H})$. Moreover, for $X_{t},Y_{t}\in[0,T]\times\mathcal{B}(\mathcal{H})$, we define the inner product:
$\langle X_{t},Y_{t}\rangle_{2,\lambda\otimes\sigma}:=\frac{1}{T}\int_{0}^{T}\langle X_{t},Y_{t}\rangle_{\sigma,1/2}\,\mathrm{d}t\,,$
where $\langle\cdot,\cdot\rangle_{\sigma,1/2}$ is the KMS inner product. For $Y_{t}=X_{t}$, we write $\lVert X_{t}\rVert_{2,\lambda\otimes\sigma}^{2}=\langle X_{t},X_{t}\rangle_{2,\lambda\otimes\sigma}$. This allows us to introduce the Sobolev space of time-dependent operators: letting $\mathfrak{S}$ be a subspace of $\mathcal{B}(\mathcal{H})$, for an integer $k\geq 0$,
$H^{k}_{\lambda\otimes\sigma}([0,T];\mathfrak{S}):=\left\{X_{t}\in[0,T]\times\mathfrak{S}\,;\ \lVert\partial_{t}^{j}X_{t}\rVert_{2,\lambda\otimes\sigma}<\infty\,,\quad j=0,1,\ldots,k\right\}\,.$
We now present the main result of this section.
Theorem 3.1.
Under A, let $\lambda_{D}$ and $s_{H}$ be constants introduced in Eqs. 3.3 and 3.7. For a given $X_{0}\in\mathcal{B}(\mathcal{H})$, we denote $X_{t}:=\mathcal{P}_{t}X_{0}$. It holds that for any $T>0$, the time average of $\lVert X_{t}\rVert_{\sigma,1/2}$ over the interval $[t,t+T]$ exponentially decays in $t$:
(3.10) $\frac{1}{T}\int_{t}^{t+T}\lVert X_{\tau}-\langle X_{\tau}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau\leq e^{-\nu t}\frac{1}{T}\int_{0}^{T}\lVert X_{\tau}-\langle X_{\tau}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau\,,$
with convergence rate:
(3.11) $\nu=\frac{\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta=0}^{2}}\,.$
Here constants $C_{1,T}$ and $C_{2,T,\beta=0}$ are defined in Theorem 3.2 below. In addition, we have
(3.12) $\lVert X_{t}-\langle X_{t}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\leq C_{T}e^{-\nu t}\lVert X_{0}-\langle X_{0}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\,,\quad\forall\,t\geq 0\,,$
where
$C_{T}:=e^{\nu T}\leq e^{\mathcal{O}\left(\frac{1}{\lambda_{D}T+(\lambda_{D}T)^{-1}}\right)}=\mathcal{O}(1)\,.$
Note that both $\nu$ and $C_{T}$ depend on the parameter $T$, which can be chosen to optimize the decay estimate (see Section 3.2). The proof follows from a quantum space-time Poincaré inequality established in Theorem 3.2 below and the standard energy estimate. For ease of readability, we shall defer all the proofs and some related technical lemmas to Section 3.3.
Theorem 3.2.
Let $s_{H}$ be the constant given in Eq. 3.7 under A. It holds that for any constant $\beta>0$ and $X_{t}\in L^{2}_{\lambda\otimes\sigma}([0,T];\mathcal{B}(\mathcal{H}))$,
$\lVert X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}\leq C_{1,T}\lVert(1-\Pi_{0})X_{t}\rVert_{2,\lambda\otimes\sigma}+C_{2,T,\beta}\lVert(\beta-\mathcal{L}^{D})^{-1/2}(-\partial_{t}+\mathcal{L}^{H})X_{t}\rVert_{2,\lambda\otimes\sigma}\,,$
where
(3.13) $C_{1,T}:=\sqrt{2}+14+\frac{12\sqrt{5}}{s_{H}T}+\frac{24}{(s_{H}T)^{2}}+\sqrt{2}\max\left\{\frac{6}{s_{H}^{2}T}+\frac{3}{s_{H}},\frac{T}{\pi}\right\}\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}\,,$
and
(3.14) $C_{2,T,\beta}:=\sqrt{2}\left\lVert(\beta-\mathcal{L}^{D})^{1/2}\right\rVert_{2\to 2}\left(\max\left\{\frac{6}{s_{H}^{2}T}+\frac{3}{s_{H}},\frac{T}{\pi}\right\}+\max\left\{\frac{1}{s_{H}},\frac{T}{\pi}\right\}\right)\,.$
Remark 7.
The blow-up of constants $C_{1,T}$ and $C_{2,T,\beta}$ in (3.13) and (3.14) as $T\to 0$ is expected from the hypocoercivity of the Lindbladian (3.1). Indeed, if there exists some constant $C>0$ such that $\limsup_{T\to 0}C_{1,T}+C_{2,T,\beta}\leq C$. Then from (3.10), we can readily derive, by letting $T\to 0$,
$\lVert X_{t}-\langle X_{t}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\leq e^{-\nu t}\lVert X_{0}-\langle X_{0}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\,,\quad\forall\,t\geq 0\,,$
for some $\nu>0$, that is, (3.12) holds with prefactor $C_{T}=1$. This means that the model is coercive, equivalently, $\dim\ker(\mathcal{L^{D}})=1$, which contradicts with A.
3.2. Scaling of the convergence rate
We next discuss the scaling of the exponential convergence rate $\nu$ in (3.12) with respect to various model parameters. This allows us to optimally select the parameters to enhance the dissipation of QMS.
Optimizing time interval length $T$. Recall that in the exponential decay (3.12), both the multiplicative constant $C_{T}$ and the convergence rate $\nu$ depend on the choice of $T$. By Eqs. 3.13 and 3.14, it holds that $C_{1,T},C_{2,T,\beta}\to+\infty$, if we set $T\to 0$ or $T\to+\infty$ (see also Remark 7). Such choice is not optimal, as in this case, the convergence rate estimate (3.11) reduces to the trivial one $\nu=0$, and Eq. 3.12 is simply the contraction: $\lVert X_{t}-\langle X_{t}\rangle_{\sigma}\rVert_{\sigma,1/2}\leq\lVert X_{0}-\langle X_{0}\rangle_{\sigma}\rVert_{\sigma,1/2}$. Thus the optimum should be achieved somewhere intermediate.
The optimal $T$ for maximizing the rate estimate could be formally written as
$T=\arg\max_{T}\frac{\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta=0}^{2}}\,.$
However, it is hard to derive its explicit formula. We will consider a few asymptotic regimes here. When $s_{H}\to 0$, using the ansatz $T=\Theta(s_{H}^{t})$ with $t\in\mathbb{R}$, we have that if $t\geq-1$,
$C_{1,T}=\Theta\left(\frac{1}{s_{H}^{2+2t}}+\frac{1}{s_{H}^{2+t}}\right)\,,\quad C_{2,T,\beta=0}=\Theta\left(\frac{1}{s_{H}^{2+t}}\right)\,,$
and if $t<-1$,
$C_{1,T}=C_{2,T,\beta=0}=\Theta\left(s_{H}^{t}\right)\,,$
which implies
$\frac{\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta=0}^{2}}=\begin{dcases}\Theta(s_{H}^{-2t})&\text{if $t\leq-1$}\,,\\ \Theta(s_{H}^{4+2t})&\text{if $0\geq t\geq-1$}\,,\\ \Theta(s_{H}^{4+4t})&\text{if $t\geq 0$}\,.\end{dcases}$
Similarly, when $s_{H}=\Theta(\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}})\to+\infty$, we have that if $t\geq-1$, then $C_{1,T}=\Theta(s_{H}^{1+t})$ and $C_{2,T,\beta=0}=\Theta(s_{H}^{t})$, while if $t<-1$, then $C_{1,T}=\Theta(s_{H}^{-2t-2})$ and $C_{2,T,\beta=0}=\Theta(s_{H}^{-t-2})$. It follows that
$\frac{\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta=0}^{2}}=\begin{dcases}\Theta(s_{H}^{4+4t})&\text{if $t\leq-1$}\,,\\ \Theta(s_{H}^{-2-2t})&\text{if $0\geq t\geq-1$}\,,\\ \Theta(s_{H}^{-2t})&\text{if $t\geq 0$}\,.\end{dcases}$
This suggests that the optimal $T$ is proportional to $s_{H}^{-1}$ (at least when $s_{H}\ll 1$ and $s_{H}\gg 1$). Therefore, we set $T=3/s_{H}$, and a direct computation gives
$C_{1,T}\leq 28+\frac{5\sqrt{2}}{s_{H}}\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}\,,\quad C_{2,T,\beta=0}=\big{\|}(\mathcal{L}^{D})^{1/2}\big{\|}_{2\to 2}\frac{6\sqrt{2}}{s_{H}}\,.$
Then, Theorem 3.1 can be simplified as follows.
Corollary 3.3.
Under A, let $\lambda_{D}$ and $s_{H}$ be constants in Eq. 3.8. For a $X_{0}\in\mathcal{B}(\mathcal{H})$, we denote $X_{t}:=\mathcal{P}_{t}X_{0}$. Then, it holds that
(3.15) $\lVert X_{t}-\langle X_{t}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\leq C_{T}e^{-\nu t}\lVert X_{0}-\langle X_{0}\rangle_{\sigma}\rVert_{\sigma,1/2}^{2}\,,\quad\forall\,t\geq 0\,,$
with the multiplicative constant $C_{T}=e^{3\nu/s_{H}}=\mathcal{O}(1)$ and the convergence rate:
(3.16) $\nu$ $=\frac{\lambda_{D}s_{H}^{2}}{(28s_{H}+5\sqrt{2}\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}})^{2}+72\lambda_{D}\big{\|}\mathcal{L}^{D}\big{\|}_{2\to 2}}$
=$=\begin{dcases}\Theta(\lambda_{D})&\text{if $s_{H}=\Theta(\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}})\gg 1$}\,,\\ \Theta\left(\frac{\lambda_{D}s_{H}^{2}}{\lVert\mathcal{L}^{H}\Pi_{+}\rVert_{\mathfrak{H}\to\mathfrak{H}}^{2}+\lambda_{D}\|\mathcal{L}^{D}\|_{2\to 2}}\right)&\text{if $s_{H}\ll 1$}\,.\end{dcases}$
Remark 8.
The lower bound estimate (3.16) for the rate $\nu$ is monotonically decreasing in $\lVert\mathcal{L}^{D}\rVert_{2\to 2}$ and $\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}$, and increasing in $\lambda_{D}$ and $s_{H}$. Thus, for a given concrete model, we need to further bound $\lambda_{D}$ and $s_{H}$ from below and $\lVert\mathcal{L}^{D}\rVert_{2\to 2}$ and $\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}$ from above, which is generally difficult and can only be considered case by case. However, a basic upper bound for $\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}$ can be given by $\lambda_{\max}(H)-\lambda_{\min}(H)$, which we will use in Section 5 where a few examples are considered. Here $\lambda_{\max/\min}(H)$ denotes the maximal/minimal eigenvalue of $H$.
Choice of coupling parameter $\alpha$. We proceed to discuss the scaling of the convergence rate $\nu$ in $\alpha$. For clarity, we will consider the Lindbladian $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ in (2.5), and suppose that A holds and operators $\mathcal{L}^{H}$ and $\mathcal{L}^{D}$ satisfy inequalities (3.3) and (3.7) with constants $\lambda_{D}$ and $s_{H}$. Note that Theorem 3.1 directly applies to the QMS generated by $\mathcal{L}_{\alpha}$.
It is convenient to introduce the rescaled Lindbladian ~$\widetilde{\mathcal{L}}_{\gamma}=\mathcal{L}^{H}+\gamma\mathcal{L}^{D}$ with $\gamma=\alpha^{-1}$ that satisfies
$e^{t\mathcal{L}_{\alpha}}=e^{\alpha t\widetilde{\mathcal{L}}_{\gamma}}\,.$
It follows that if the exponential decay (3.12) holds for the flow $X_{t}=e^{t\widetilde{\mathcal{L}}_{\gamma}}X_{0}$ with the prefactor $C_{T}$ and the convergence rate $\nu(\widetilde{\mathcal{L}}_{\gamma})$, then $X_{t}=e^{t\mathcal{L}_{\alpha}}X_{0}$ would satisfy (3.12) with the same $C_{T}$ and rate $\nu(\mathcal{L}_{\alpha})=\alpha\nu(\widetilde{\mathcal{L}}_{\gamma})$. We apply Theorem 3.1 to ~$\widetilde{\mathcal{L}}_{\gamma}$ and find
$\nu(\widetilde{\mathcal{L}}_{\gamma})\geq\frac{\gamma\lambda_{D}}{C_{1,T}^{2}+\gamma^{2}\lambda_{D}C_{2,T,\beta=0}^{2}}\,,$
which is maximized at
$\gamma=\frac{C_{1,T}}{\sqrt{\lambda_{D}}C_{2,T,\beta=0}}\underset{\begin{subarray}{c}s_{H}\ll 1\\ T=3/s_{H}\end{subarray}}{=}\Theta\left(\frac{\lVert\mathcal{L}^{H}\Pi_{+}\rVert_{\mathfrak{H}\to\mathfrak{H}}}{\sqrt{\lambda_{D}\|\mathcal{L}^{D}\|_{2\to 2}}}\right)\,.$
Then, we readily obtain
(3.17) $\nu(\mathcal{L}_{\alpha})=\alpha\nu(\widetilde{\mathcal{L}}_{\gamma})\geq\frac{\lambda_{D}}{C_{1,T}^{2}+\alpha^{-2}\lambda_{D}C_{2,T,\beta=0}^{2}}\nearrow\frac{\lambda_{D}}{C_{1,T}^{2}}\quad\text{as}\quad\alpha\to\infty\,,$
i.e., the lower bound estimate of $\nu$ is monotonically increasing when $\alpha\to\infty$. In this case, we can further derive, when $T=3/s_{H}$,
$\liminf_{\alpha\to\infty}\nu(\mathcal{L}_{\alpha})\geq\frac{\lambda_{D}}{C_{1,T}^{2}}\underset{s_{H}\ll 1}{=}\Theta\left(\frac{\lambda_{D}s_{H}^{2}}{\lVert\mathcal{L}^{H}\Pi_{+}\rVert^{2}_{\mathfrak{H}\to\mathfrak{H}}}\right)\,.$
We emphasize that the scaling behaviors of $\nu(\mathcal{L}_{\alpha})$ and $\nu(\widetilde{\mathcal{L}}_{\gamma})$ are quite different. In Section 4, we shall analyze the spectral gap $\lambda(\mathcal{L}_{\alpha})$ in the asymptotic regime $\alpha\to\infty$, which provides a more accurate convergence rate estimation for large $\alpha$. Some further straightforward estimation, along with the above discussions and Corollary 3.3, gives the following results.
Corollary 3.4.
Under A, let $\lambda_{D}$ and $s_{H}$ be constants in Eq. 3.8 with $s_{H}\ll 1$.
  • For the Lindbladian ~$\widetilde{\mathcal{L}}_{\gamma}=\mathcal{L}^{H}+\gamma\mathcal{L}^{D}$ with $\gamma>0$, the convergence rate estimate (3.16) is maximized at $\gamma_{*}=\Theta\left(\frac{\lVert\mathcal{L}^{H}\Pi_{+}\rVert_{\mathfrak{H}\to\mathfrak{H}}}{\sqrt{\lambda_{D}\|\mathcal{L}^{D}\|_{2\to 2}}}\right)$, and there holds
    $\max_{\gamma>0}\nu(\widetilde{\mathcal{L}}_{\gamma})\geq\Theta\left(\frac{s_{H}^{2}\sqrt{\lambda_{D}}}{\lVert\mathcal{L}^{H}\Pi_{+}\rVert_{\mathfrak{H}\to\mathfrak{H}}\sqrt{\|\mathcal{L}^{D}\|_{2\to 2}}}\right)\,.$
  • For the Lindbladian $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ with $\alpha>0$, the convergence rate estimate (3.16) is increasing in $\alpha$, and there holds
    $\liminf_{\alpha\to\infty}\nu(\mathcal{L}_{\alpha})\geq\Theta\left(\frac{s_{H}^{2}\lambda_{D}}{\lVert\mathcal{L}^{H}\Pi_{+}\rVert^{2}_{\mathfrak{H}\to\mathfrak{H}}}\right)\,.$
Comparison with [fang2024mixing] via DMS approach. We briefly recall the results in [fang2024mixing] that is motivated by [dolbeault2015hypocoercivity], as a comparison. We still assume A for a Lindbladian $\mathcal{L}=\mathcal{L}^{H}+\mathcal{L}^{D}$ and let constants $s_{H}$ and $\lambda_{D}$ be given as in (3.8). For some $\eta>0$, we define the operator:
$\mathcal{A}:=\left(\eta+(\mathcal{L}^{H}\Pi_{0})^{\star}(\mathcal{L}^{H}\Pi_{0})\right)^{-1}(\mathcal{L}^{H}\Pi_{0})^{\star}\,,$
and the Lyapunov functional for $\epsilon\in(0,1)$:
$\mathsf{L}(X)=\frac{1}{2}\lVert X\rVert_{\sigma,1/2}^{2}-\epsilon\,\mathfrak{Re}\langle\mathcal{A}X,X\rangle_{\sigma,1/2}\,.$
Let the constant $C_{M}(\eta)$ be given by
$\lVert\mathcal{A}\mathcal{L}^{H}\Pi_{+}X\rVert_{\sigma,1/2}+\lVert\mathcal{A}\mathcal{L}^{D}X\rVert_{\sigma,1/2}\leq C_{M}(\eta)\lVert\Pi_{+}X\rVert_{\sigma,1/2}\,.$
Then, we have the exponential decay estimate [fang2024mixing]*Theorem 1:
$\lVert e^{t\mathcal{L}}X\rVert_{\sigma,1/2}\leq Ce^{-\nu t}\lVert X\rVert_{\sigma,1/2},,\quad\text{for}\ X\in\mathfrak{H}\,,$
where
(3.18) $C=\left(\frac{1+\epsilon}{1-\epsilon}\right)^{1/2}\,,\quad\nu=\min\left\{\frac{1}{4}\frac{\lambda_{D}}{1+\epsilon},\frac{1}{3}\frac{\epsilon s_{H}^{2}}{(1+\epsilon)(\eta+s_{H}^{2})}\right\},$
with
(3.19) $\epsilon=\frac{1}{2}\min\left\{\frac{\lambda_{D}s_{H}^{2}}{(\eta+s_{H}^{2})(1+C_{M}(\eta))^{2}},1\right\}\,.$
To compare with our results, we let $\lambda_{D},s_{H}\ll 1$, and note, by using [fang2024mixing]*Lemma 2,
$C_{M}(\eta)\leq\frac{1}{2\sqrt{\eta}}\left(\lVert\mathcal{L}^{H}\Pi_{+}\rVert_{\mathfrak{H}\to\mathfrak{H}}+\lVert\mathcal{L}^{D}\rVert_{2\to 2}\right)\,.$
It follows that the optimal $\eta$ in this regime is $\eta=\Theta(s_{H}^{2})$, and there holds
$C=\Theta(1)\,,\quad\nu=\Theta\left(\frac{\lambda_{D}s_{H}^{2}}{\lVert\mathcal{L}^{H}\Pi_{+}\rVert_{\mathfrak{H}\to\mathfrak{H}}+\lVert\mathcal{L}^{D}\rVert_{2\to 2}}\right)\,,$
which matches the rate estimate in Corollary 3.3 in the sense that both give $\nu=\Theta(\lambda_{D}s_{H}^{2})$. However, beyond the asymptotic regime $\lambda_{D},s_{H}\ll 1$, accurate estimations for $\lambda_{D}$, $s_{H}$, and $C_{M}(\eta)$ are necessary to find the minimum in Eqs. 3.18 and 3.19. Moreover, it is challenging to determine the optimal $\gamma$ and $\alpha$ for the convergence rate estimation in (3.18), as in Corollary 3.4. One advantage of our results is that they do not depend on the twisted norm, and their scaling with respect to the model parameters is more explicit. It would be interesting to conduct a detailed case study for various physical models to understand the sharpness of our decay rate estimate compared to that in [fang2024mixing].
3.3. Proofs and Lemmas
This section is devoted to the proofs of Theorem 3.1 (exponential decay of time-averaged $L^{2}$-distance) and Theorem 3.2 (quantum space-time Poincaré inequality). For this, we need some auxiliary technical lemmas, in particular, the key Lemma 3.7, where we consider a divergence-type equation, whose solutions provide some crucial test functions used for the proof of Theorem 3.2. We start with lifting Eq. 3.7 to the space of time-dependent operators, which follows from a standard tensorization technique. For ease of exposition, we define the subspace of $\mathcal{B}(\mathcal{H})$:
[…]
Lemma 3.5.
[…]
(3.20) $\lVert X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert^{2}_{2,\lambda\otimes\sigma}\leq\max\left\{\frac{1}{s_{H}^{2}},\frac{T^{2}}{\pi^{2}}\right\}\left(\left\lVert\mathcal{L}_{+0}^{H}X_{t}\right\rVert_{2,\lambda\otimes\sigma}^{2}+\big{\|}\partial_{t}X_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}\right).$
Proof.
We first note
$\lVert X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert^{2}_{2,\lambda\otimes\sigma}=\lVert X_{t}-\langle X_{t}\rangle_{\lambda}\rVert^{2}_{2,\lambda\otimes\sigma}+\lVert\langle X_{t}\rangle_{\lambda}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
thanks to, for any $X_{t}\in L^{2}_{\lambda\otimes\sigma}([0,T];\mathcal{B}(\mathcal{H}))$,
$\langle X_{t}-\langle X_{t}\rangle_{\lambda},\langle X_{t}\rangle_{\lambda}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rangle_{2,\lambda\otimes\sigma}=0\,.$
[…]
$\lVert\langle X_{t}\rangle_{\lambda}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}^{2}\leq\frac{1}{s^{2}_{H}}\lVert\langle\mathcal{L}_{+0}^{H}X_{t}\rangle_{\lambda}\rVert_{\sigma,1/2}^{2}\leq\frac{1}{s^{2}_{H}}\lVert\mathcal{L}_{+0}^{H}X_{t}\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
where the last inequality follows from the convexity of $\lVert\cdot\rVert_{\sigma,1/2}^{2}$. Moreover, by the standard Poincaré inequality for $H^{1}$-functions [evans2022partial]:
$\left\|f-\int_{0}^{T}f(s)\,\mathrm{d}\lambda\right\|_{L^{2}([0,T])}\leq\frac{T}{\pi}\lVert f^{\prime}\rVert_{L^{2}([0,T])}\,,$
there holds
$\lVert X_{t}-\langle X_{t}\rangle_{\lambda}\rVert^{2}_{2,\lambda\otimes\sigma}$ $=\int_{0}^{T}\lVert X_{t}-\langle X_{t}\rangle_{\lambda}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\lambda$
$\leq\frac{T^{2}}{\pi^{2}}\int_{0}^{T}\lVert\partial_{t}X_{t}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\lambda\leq\frac{T^{2}}{\pi^{2}}\lVert\partial_{t}X_{t}\rVert_{2,\lambda\otimes\sigma}^{2}\,.$
Then, we have
$\lVert X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert^{2}_{2,\lambda\otimes\sigma}\leq\frac{1}{s^{2}_{H}}\lVert\mathcal{L}_{+0}^{H}X_{t}\rVert_{2,\lambda\otimes\sigma}^{2}+\frac{T^{2}}{\pi^{2}}\lVert\partial_{t}X_{t}\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
which gives the desired (3.20). ∎
We consider the following abstract elliptic operator:
(3.21) $\mathfrak{L}^{H}=-\partial_{tt}+\big{(}\mathcal{L}_{+0}^{H}\big{)}^{\star}\mathcal{L}_{+0}^{H}\,,$
and the space:
[…]
The well-posedness of the operator $\mathfrak{L}^{H}$ is established in the following lemma.
Lemma 3.6.
[…]
(3.22) $\mathfrak{L}^{H}X_{t}=F_{t}\,,\quad\partial_{t}X_{t}|_{t=0,T}=0\,.$
Then there is a unique weak solution $X_{t}\in\mathcal{V}$ such that
[…]
(3.24) $\lVert\partial_{t}X_{t}\rVert^{2}_{2,\lambda\otimes\sigma}+\lVert\mathcal{L}_{+0}^{H}X_{t}\rVert_{2,\lambda\otimes\sigma}^{2}\leq\max\left\{\frac{1}{s^{2}_{H}},\frac{T^{2}}{\pi^{2}}\right\}\lVert F_{t}\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
and
(3.25) $\big{\|}\partial_{tt}X_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}+\left\lVert(\mathcal{L}_{+0}^{H})^{\star}\mathcal{L}_{+0}^{H}X_{t}\right\rVert_{2,\lambda\otimes\sigma}^{2}+2\big{\|}\partial_{t}\mathcal{L}_{+0}^{H}X_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}=\big{\|}\mathfrak{L}^{H}X_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}=\big{\|}F_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}\,.$
Proof.
We consider the bilinear form on $\mathcal{V}$ associated with the equation (3.22):
$\mathcal{Q}(Y_{t},X_{t})=\big{\langle}\partial_{t}Y_{t},\partial_{t}X_{t}\big{\rangle}_{2,\lambda\otimes\sigma}+\big{\langle}\mathcal{L}_{+0}^{H}Y_{t},\mathcal{L}_{+0}^{H}X_{t}\big{\rangle}_{2,\lambda\otimes\sigma}\,.$
One can easily see that $\mathcal{Q}(\cdot,\cdot)$ is an inner product on $\mathcal{V}$. In fact, if $\mathcal{Q}(X,X)=0$, we have $\partial_{t}X_{t}=0$ and $\mathcal{L}_{+0}^{H}X=0$. The latter condition $\mathcal{L}_{+0}^{H}X=0$ gives that for a.e. $t\in[0,T]$, $X_{t}$ is proportional to the identity ${\bf 1}$. Then the former one $\partial_{t}X_{t}=0$, along with $\langle X_{t}\rangle_{\lambda\otimes\sigma}=0$, implies $X_{t}=0$. Further, Lemma 3.5 shows that the bilinear form $\mathcal{Q}(\cdot,\cdot)$ is coercive on $\mathcal{V}$: for some $C>0$,
$\mathcal{Q}(X_{t},X_{t})\geq C\left(\lVert X_{t}\rVert^{2}_{2,\lambda\otimes\sigma}+\lVert\partial_{t}X_{t}\rVert^{2}_{2,\lambda\otimes\sigma}\right),\quad\forall\,X\in\mathcal{V}\,.$
The existence of the weak solution to (3.22) is guaranteed by the Lax-Milgram theorem with the estimate (3.24). Finally, note that $\mathfrak{L}^{H}X_{t}=F_{t}$ is equivalent to
$-\partial_{tt}X_{t}=F_{t}-\left(\mathcal{L}_{+0}^{H}\right)^{\star}\mathcal{L}_{+0}^{H}X_{t}\,,$
[…]
$\big{\|}\mathfrak{L}^{H}X_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}$ $=\big{\langle}(-\partial_{tt}+(\mathcal{L}_{+0}^{H})^{\star}\mathcal{L}_{+0}^{H})X_{t},(-\partial_{tt}+(\mathcal{L}_{+0}^{H})^{\star}\mathcal{L}_{+0}^{H})X_{t}\big{\rangle}_{2,\lambda\otimes\sigma}$
$=\big{\|}\partial_{tt}X_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}+\left\lVert(\mathcal{L}_{+0}^{H})^{\star}\mathcal{L}_{+0}^{H}X_{t}\right\rVert_{2,\lambda\otimes\sigma}^{2}+2\,\mathfrak{Re}\left\langle-\partial_{tt}X_{t},(\mathcal{L}_{+0}^{H})^{\star}\mathcal{L}_{+0}^{H}X_{t}\right\rangle_{2,\lambda\otimes\sigma}\,,$
and by the commutativity between $\partial_{t}$ and $\mathcal{L}_{+0}^{H}$,
$\left\langle-\partial_{tt}X_{t},(\mathcal{L}_{+0}^{H})^{\star}\mathcal{L}_{+0}^{H}X_{t}\right\rangle_{2,\lambda\otimes\sigma}=\left\langle\partial_{t}\mathcal{L}_{+0}^{H}X_{t},\partial_{t}\mathcal{L}_{+0}^{H}X_{t}\right\rangle_{2,\lambda\otimes\sigma}\,.\qed$
The next technical lemma is essential and deals with a divergence-type equation (3.26) associated with the elliptic one (3.22), which is inspired by the arguments in [cao2023explicit]*Lemma 2.6 and [brigati2023construct]*Section 6 for the underdamped Langevin dynamics.
Lemma 3.7.
[…]
(3.26) $-\partial_{t}Y^{(0)}_{t}+\left(\mathcal{L}_{+0}^{H}\right)^{\star}\ Y_{t}^{(1)}=F_{t}\,,\quad\text{$Y^{(j)}_{t}|_{t=0,T}=0$ for $j=0,1$}\,.$
In addition, we have
(3.27) $\big{\|}Y^{(0)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\leq\sqrt{2}\max\left\{\frac{1}{s_{H}},\frac{T}{\pi}\right\}\left\lVert F_{t}\right\rVert_{2,\lambda\otimes\sigma}\,,$
$\big{\|}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\leq\sqrt{2}\max\left\{\frac{6}{s_{H}^{2}T}+\frac{3}{s_{H}},\frac{T}{\pi}\right\}\left\lVert F_{t}\right\rVert_{2,\lambda\otimes\sigma}\,,$
and
(3.28) $\big{\|}\partial_{t}Y^{(0)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\leq\sqrt{26}\left(1+\frac{1}{\sqrt{26}}+\frac{2}{s_{H}T}\right)\left\lVert F_{t}\right\rVert_{2,\lambda\otimes\sigma}\,,$
$\big{\|}\partial_{t}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\leq 12\left(\frac{13}{12}+\frac{\sqrt{5}}{s_{H}T}+\frac{2}{(s_{H}T)^{2}}\right)\left\lVert F_{t}\right\rVert_{2,\lambda\otimes\sigma}\,.$
Proof.
[…]
(3.29) $F_{t}=F^{h}_{t}+F^{\perp}_{t}\quad\text{with}\quad F^{h}_{t}\in\mathbb{H}\,,\ F^{\perp}_{t}\perp\mathbb{H}\,,$
with respect to the inner product $\langle\cdot,\cdot\rangle_{2,\lambda\otimes\sigma}$. Since a constant operator $X_{t}=c{\bf 1}$ with $c\in\mathbb{C}$ belongs to $\mathbb{H}$, we have $\langle F^{\perp}_{t}\rangle_{\lambda\otimes\sigma}=0$ and thus $\langle F^{h}_{t}\rangle_{\lambda\otimes\sigma}=\langle F-F^{\perp}_{t}\rangle_{\lambda\otimes\sigma}=0$. We shall construct the solution to the equation (3.26) with source $F^{h}_{t}$ or $F^{\perp}_{t}$ separately.
Construction of solution for $F^{\perp}$. We first consider the easier case of $F^{\perp}_{t}$. Let $X_{t}$ be the solution to the variational equation (3.23) with source $F^{\perp}_{t}$. We take the test function $Z_{t}\in\mathbb{H}$ and obtain
$0=\left\langle Z_{t},F_{t}^{\perp}\right\rangle_{2,\lambda\otimes\sigma}=\left\langle\mathfrak{L}^{H}Z_{t},X_{t}\right\rangle_{2,\lambda\otimes\sigma}+\left\langle\partial_{t}Z_{t},X_{t}\right\rangle_{\sigma,1/2}|_{t=T}-\left\langle\partial_{t}Z_{t},X_{t}\right\rangle_{\sigma,1/2}|_{t=0}\,,$
using integration by parts, thanks to the $H^{2}$-regularity of $Z_{t}$ in $t$. Noting that $\mathfrak{L}^{H}Z_{t}=0$ and $\partial_{t}Z_{t}$ is arbitrary at $t=0,T$ (see also the decomposition (3.32) below for an operator in $\mathbb{H}$), we can conclude $X_{t}|_{t=0,T}=0$. We then define
$Y^{(0)}_{t}=\partial_{t}X_{t}\,,\quad Y^{(1)}_{t}=\mathcal{L}_{+0}^{H}X_{t}\,,$
with the following estimate by Lemma 3.6,
(3.30) $\big{\|}Y^{(0)}_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}+\big{\|}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}\leq\max\left\{\frac{1}{s^{2}_{H}},\frac{T^{2}}{\pi^{2}}\right\}\left\lVert F^{\perp}_{t}\right\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
and
(3.31) $\big{\|}\partial_{t}Y^{(0)}_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}+\big{\|}\partial_{t}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}\leq\left\lVert\mathfrak{L}^{H}X_{t}\right\rVert_{2,\lambda\otimes\sigma}^{2}\leq\left\lVert F^{\perp}_{t}\right\rVert_{2,\lambda\otimes\sigma}^{2}\,.$
Construction of solution for $F^{h}$. We proceed to consider the case of $F^{h}_{t}$. For this, recall that $\left(\mathcal{L}_{+0}^{H}\right)^{\star}\mathcal{L}_{+0}^{H}=\Pi_{0}\left(\mathcal{L}^{H}\right)^{2}\Pi_{0}$ is a self-adjoint operator on $\mathfrak{H}_{0}$. We assume that its orthonormal eigenvectors $\{W_{\mu}\}$ are given by
$\left(\mathcal{L}_{+0}^{H}\right)^{\star}\mathcal{L}_{+0}^{H}W_{\mu}=\mu^{2}W_{\mu}\,,$
where $\mu^{2}\geq s_{H}^{2}>0$ by Eq. 3.7. With the help of $\{W_{\mu}\}$, we have the following decomposition:
(3.32) $F^{h}_{t}=f_{0}(t)+\sum_{\mu}f_{\mu}(t)W_{\mu}\,,$
with $\langle f_{0}(t)\rangle_{\lambda}=0$. Then the condition $\mathfrak{L}^{H}F_{t}^{h}=0$ gives
$0=-f_{0}^{\prime\prime}(t)+\sum_{\mu}\left(-f_{\mu}^{\prime\prime}+\mu^{2}f_{\mu}(t)\right)W_{\mu}\,,$
further implying
$f_{0}^{\prime\prime}(t)=0\qquad\text{and}\qquad f_{\mu}^{\prime\prime}-\mu^{2}f_{\mu}(t)=0\,.$
An elementary calculation yields $f_{0}(t)=c(t-\frac{T}{2})$ for some constant $c\in\mathbb{C}$, and that $f_{\mu}(t)$ can be decomposed as: for some $c_{+}^{\mu},c_{-}^{\mu}\in\mathbb{C}$,
(3.33) $f_{\mu}(t)=c_{+}^{\mu}e^{-\mu t}+c_{-}^{\mu}e^{-\mu(T-t)}\,,$
where coefficients $c_{\pm}^{\mu}$ are computed below. Due to linearity, to construct the solution to Eq. 3.26 for $F^{h}_{t}$, it suffices to construct for each term in the decomposition (3.32) and add them together.
Computation of coefficients $c_{\pm}^{\mu}$. We now calculate the constants $c_{\pm}^{\mu}$ in (3.33). For convenience, we define constants:
(3.34) $v_{T}:=\lVert e^{-\mu t}\rVert_{2,\lambda}^{2}=\lVert e^{-\mu(T-t)}\rVert_{2,\lambda}^{2}=\int_{0}^{T}e^{-2\mu t}\,\mathrm{d}\lambda(t)=\frac{1-\ell^{2}}{2\mu T}\,,\quad\ell:=e^{-\mu T}\,,$
and functions ${\bf e}_{\mu,\pm}$ as linear combinations of $e^{-\mu t}$ and $e^{-\mu(T-t)}$:
${\bf e}_{\mu,+}(t):=e^{-\mu t}-\frac{\langle e^{-\mu t},e^{-\mu(T-t)}\rangle_{2,\lambda}}{\lVert e^{-\mu(T-t)}\rVert_{2,\lambda}^{2}}e^{-\mu(T-t)}=e^{-\mu t}-v_{T}^{-1}e^{-\mu(2T-t)}\,,$
and
${\bf e}_{\mu,-}(t):=e^{-\mu(T-t)}-\frac{\langle e^{-\mu t},e^{-\mu(T-t)}\rangle_{2,\lambda}}{\lVert e^{-\mu t}\rVert_{2,\lambda}^{2}}e^{-\mu t}=e^{-\mu(T-t)}-v_{T}^{-1}e^{-\mu(t+T)}\,.$
Taking the inner product of Eq. 3.33 with $e^{-\mu t}$ and $e^{-\mu(T-t)}$ and solving the resulting linear equations in $c_{\pm}^{\mu}\in\mathbb{C}$, we find
$c_{+}^{\mu}:=\frac{\langle f_{\mu},{\bf e}_{\mu,+}\rangle_{2\,\lambda}}{\langle{\bf e}_{\mu,+},e^{-\mu t}\rangle_{2,\lambda}}\,,\quad c_{-}^{\mu}:=\frac{\langle f_{\mu},{\bf e}_{\mu,-}\rangle_{2\,\lambda}}{\langle{\bf e}_{\mu,-},e^{-\mu(T-t)}\rangle_{2,\lambda}}\,.$
Before we proceed, we estimate
(3.35) $\lVert c_{+}^{\mu}e^{-\mu t}\rVert_{2,\lambda}^{2}=|c_{+}^{\mu}|^{2}v_{T}\leq\left(\frac{\lVert{\bf e}_{\mu,+}\rVert_{2,\lambda}}{\langle{\bf e}_{\mu,+},e^{-\mu t}\rangle_{2,\lambda}}\right)^{2}\lVert f_{\mu}\rVert_{2,\lambda}^{2}v_{T}=\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\lVert f_{\mu}\rVert_{2,\lambda}^{2}\,,$
by a direct computation:
$\lVert{\bf e}_{\mu,+}\rVert_{2,\lambda}^{2}=\langle{\bf e}_{\mu,+},e^{-\mu t}\rangle_{2,\lambda}=v_{T}-v_{T}^{-1}\ell^{2}\,.$
It is also easy to see that $v_{T}^{2}/(v_{T}^{2}-\ell^{2})$ is monotonically decreasing in $\mu T$ and
$\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\to\begin{dcases}1\,,&{\rm as}\quad\mu T\to\infty\,,\\ \frac{3}{(\mu T)^{2}}(1+\mathcal{O}(\mu T))\,,&{\rm as}\quad\mu T\to 0\,.\\ \end{dcases}$
Thus, there exists a universal constant $C$ such that, for any $\mu,T>0$,
(3.36) $\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\leq\frac{4}{(\mu T)^{2}}+1\,.$
A similar estimate holds for $\lVert c_{-}^{\mu}e^{-\mu(T-t)}\rVert_{2,\lambda}$:
$\lVert c_{-}^{\mu}e^{-\mu(T-t)}\rVert_{2,\lambda}^{2}=|c_{-}^{\mu}|^{2}v_{T}\leq\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\lVert f_{\mu}\rVert_{2,\lambda}^{2}\,.$
Construction of solution for $F^{h}=f_{0}(t)$. We start with constructing the solution for $f_{0}(t)$, which is direct. We set $Y^{(0)}_{t}=-\int_{0}^{t}f_{0}(\tau)\,\mathrm{d}\tau$ and $Y^{(1)}_{t}=0$. It is easy to see, by $Y_{t}^{(0)}|_{t=0,T}=0$,
(3.37) $\big{\|}Y^{(0)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\leq\frac{T}{\pi}\lVert f_{0}\rVert_{2,\lambda\otimes\sigma}\,,$
and by definition of $Y_{t}^{(0)}$,
(3.38) $\big{\|}\partial_{t}Y^{(0)}_{t}\big{\|}_{2,\lambda\otimes\sigma}=\lVert f_{0}\rVert_{2,\lambda\otimes\sigma}\,.$
Construction of solution for $F^{h}=e^{-\mu t}W_{\mu}$. We next consider the case of $f_{\mu}(t)W_{\mu}$ with $f_{\mu}$ being of the form (3.33). By linear combination, we only need to consider $f_{\mu}(t)=e^{-\mu t}$ or $e^{-\mu(T-t)}$. We shall focus on the former case, as the latter one can be similarly handled by reversing the time. To be precise, we shall construct the solution to
$-\partial_{t}Y^{(0)}_{t}+\left(\mathcal{L}_{+0}^{H}\right)^{\star}\ Y_{t}^{(1)}=e^{-\mu t}W_{\mu}\,.$
We consider the following ansatz: $Y^{(0)}_{t}=\psi_{0,\mu}(t)W_{\mu}$ and $Y^{(1)}_{t}=\psi_{1,\mu}(t)\mathcal{L}^{H}_{+0}W_{\mu}$ for some functions $\psi_{0,\mu}$ and $\psi_{1,\mu}$ satisfying $\psi_{0,\mu}|_{t=0,T}=\psi_{1,\mu}|_{t=0,T}=0$ and the equation:
(3.39) $-\psi_{0,\mu}^{\prime}(t)+\mu^{2}\psi_{1,\mu}(t)=e^{-\mu t}\,.$
We first compute
(3.40) $\lVert e^{-\mu t}W_{\mu}\rVert_{2,\lambda\otimes\sigma}^{2}=\frac{1}{T}\int_{0}^{T}e^{-2\mu t}\,\mathrm{d}t=v_{T}\,.$
Motivated by the proof of [cao2023explicit]*Lemma 2.6, we consider the following ansatz:
(3.41) $\psi_{0,\mu}(t)=p_{0,\mu}(e^{-\mu t})\,,\quad\psi_{1,\mu}(t)=p_{1,\mu}(e^{-\mu t})\,,$
where $p_{0,\mu}$ and $p_{1,\mu}$ are polynomial functions to be constructed. Note from the equation (3.39) that (recalling $\ell=e^{-\mu T}$ in Eq. 3.34)
(3.42) $\int_{0}^{T}\psi_{1,\mu}(t)\,\mathrm{d}t=\int_{\ell}^{1}\frac{p_{1,\mu}(x)}{\mu x}\,\mathrm{d}x=\int_{0}^{T}\frac{e^{-\mu t}+\psi_{0,\mu}^{\prime}(t)}{\mu^{2}}\,\mathrm{d}t=\frac{1-\ell}{\mu^{3}}\,.$
We define
(3.43) $p_{1,\mu}(x)=\mu^{-2}g_{\mu}(x)\,,\quad g_{\mu}(x)=\frac{6}{(\ell-1)^{2}}x(1-x)(x-\ell)\,,$
which satisfies the constraint (3.42) and $p_{1,\mu}(1)=p_{1,\mu}(\ell)=0$. We then compute
(3.44) $\psi_{0,\mu}(t)=\int_{0}^{t}\mu^{2}p_{1,\mu}(e^{-\mu s})-e^{-\mu s}\,\mathrm{d}s=\frac{1}{\mu}\int_{e^{-\mu t}}^{1}\frac{g_{\mu}(x)-x}{x}\,\mathrm{d}x=p_{0,\mu}(e^{-\mu t})\,,$
and by Eq. 3.43,
(3.45) $\mu p_{0,\mu}(s)=\int_{s}^{1}\frac{6}{(\ell-1)^{2}}(1-x)(x-\ell)-1\,\mathrm{d}x=\frac{(s-1)(\ell-2s+1)(\ell-s)}{(\ell-1)^{2}}\,.$
By the construction, we have, for $x\in[\ell,1]$,
$0\leq g_{\mu}(x)\leq\frac{3}{2}x\,,\quad g_{\mu}^{\prime}(x)=\frac{6}{(\ell-1)^{2}}(-3x^{2}-\ell+2x(1+\ell))\,.$
It follows that, by Eq. 3.42,
(3.46) $\lVert\psi_{1,\mu}\rVert_{2,\lambda}^{2}=\frac{1}{T\mu^{5}}\int_{\ell}^{1}\frac{g_{\mu}(x)^{2}}{x}\,\mathrm{d}x\leq\frac{9}{4T\mu^{5}}\int_{\ell}^{1}x\,\mathrm{d}x\leq\frac{9(1-\ell^{2})}{8T\mu^{5}}=\frac{9}{4\mu^{4}}v_{T}\,,$
and
(3.47) $\lVert\psi_{1,\mu}^{\prime}\rVert_{2,\lambda}^{2}=\frac{1}{T\mu^{3}}\int_{\ell}^{1}g^{\prime}(x)^{2}x\,\mathrm{d}x$ $=\frac{36}{T\mu^{3}(\ell-1)^{4}}\int_{\ell}^{1}(-3x^{2}-\ell+x(2+2\ell))^{2}x\,\mathrm{d}x$
$=\frac{6}{5T\mu^{3}(\ell-1)^{4}}(1-\ell)^{3}(3+2\ell+2\ell^{2}+3\ell^{3})\leq\frac{12}{T\mu^{3}(1-\ell)}\,.$
Noting $\psi_{0,\mu}^{\prime}(t)=\mu^{2}\psi_{1,\mu}(t)-e^{\mu t}$ by Eq. 3.39, a straightforward computation gives
(3.48) $\lVert\psi_{0,\mu}^{\prime}\rVert_{2,\lambda}^{2}\leq 2\mu^{4}\lVert\psi_{1,\mu}\rVert_{2,\lambda}^{2}+2\lVert e^{\mu t}\rVert_{2,\lambda}^{2}\leq 2\mu^{4}\frac{9}{4\mu^{4}}v_{T}+2v_{T}=\frac{13}{2}v_{T}\,.$
Then, by Eqs. 3.44 and 3.45, we have
(3.49) $\lVert\psi_{0,\mu}\rVert_{2,\lambda}^{2}=\frac{1}{T\mu}\int_{\ell}^{1}\frac{p_{0,\mu}(x)^{2}}{x}\,\mathrm{d}x$ $=\frac{(1-\ell)^{3}}{T\mu^{3}}\int_{0}^{1}\frac{(s-1)^{2}(2s-1)^{2}s^{2}}{s+\ell(1-s)}\,\mathrm{d}s$
$\leq\frac{(1-\ell)^{3}}{T\mu^{3}}\int_{0}^{1}(s-1)^{2}(2s-1)^{2}s\,\mathrm{d}s=\frac{(1-\ell)^{2}}{30(1+\ell)\mu^{2}}v_{T}\,.$
Estimate of norms. Recalling the decomposition (3.32) for $F^{h}_{t}$, with the help of $\psi_{0,\mu}$ and $\psi_{1,\mu}$ defined above, the solution to Eq. 3.26 with source $F^{h}_{t}$ can be constructed as follows:
(3.50) $Y_{t}^{(0)}=-\int_{0}^{t}f_{0}(\tau)\,\mathrm{d}\tau+\sum_{\mu}\left(c_{+}^{\mu}\psi_{0,\mu}(t)-c_{-}^{\mu}\psi_{0,\mu}(T-t)\right)W_{\mu}\,,$
and
(3.51) $Y_{t}^{(1)}=\sum_{\mu}\left(c_{+}^{\mu}\psi_{1,\mu}(t)+c_{-}^{\mu}\psi_{1,\mu}(T-t)\right)\mathcal{L}^{H}_{+0}W_{\mu}\,.$
We now estimate, by Eq. 3.37,
$\big{\|}Y^{(0)}_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}$ $\leq\frac{T^{2}}{\pi^{2}}\lVert f_{0}\rVert_{2,\lambda}^{2}+\sum_{\mu}\lVert c_{+}^{\mu}\psi_{0,\mu}(t)-c_{-}^{\mu}\psi_{0,\mu}(T-t)\rVert_{2,\lambda}^{2}$
$\leq\frac{T^{2}}{\pi^{2}}\lVert f_{0}\rVert_{2,\lambda}^{2}+2\sum_{\mu}(|c_{+}^{\mu}|^{2}+|c_{-}^{\mu}|^{2})\lVert\psi_{0,\mu}(t)\rVert_{2,\lambda}^{2}\,,$
and similarly,
$\big{\|}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}$ $\leq\sum_{\mu}\lVert c_{+}^{\mu}\psi_{1,\mu}(t)+c_{-}^{\mu}\psi_{1,\mu}(T-t)\rVert^{2}_{2,\lambda}\mu^{2}$
$\leq 2\mu^{2}\sum_{\mu}(|c_{+}^{\mu}|^{2}+|c_{-}^{\mu}|^{2})\lVert\psi_{1,\mu}(t)\rVert_{2,\lambda}^{2}\,.$
Then, thanks to Eq. 3.49, we have, by $\mu^{2}\geq s^{2}_{H}$ and Eq. 3.35,
(3.52) $\sum_{\mu}|c_{+}^{\mu}|^{2}\lVert\psi_{0,\mu}(t)\rVert_{2,\lambda}^{2}\leq$ $\sum_{\mu}|c_{+}^{\mu}|^{2}\frac{(1-\ell)^{2}}{30(1+\ell)}\frac{1}{s^{2}_{H}}v_{T}$
$\leq$ $\sum_{\mu}\frac{(1-\ell)^{2}}{30(1+\ell)}\frac{1}{s^{2}_{H}}\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\lVert f_{\mu}\rVert_{2,\lambda}^{2}$
$\leq$ $\,\frac{1}{20}\frac{1}{s^{2}_{H}}\sum_{\mu}\lVert f_{\mu}\rVert_{2,\lambda}^{2}\,,$
for a universal constant $C>0$, where the third inequality follows from
$\frac{(1-\ell)^{2}}{1+\ell}\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\leq\frac{3}{2}\,,\quad\mu T>0\,.$
Similarly, by Eqs. 3.36 and 3.46, we have
(3.53) $\sum_{\mu}|c_{+}^{\mu}|^{2}\mu^{2}\lVert\psi_{1,\mu}(t)\rVert_{2,\lambda}^{2}\leq$ $\sum_{\mu}|c_{+}^{\mu}|^{2}\frac{9}{4s_{H}^{2}}v_{T}\leq\sum_{\mu}\frac{9}{4s^{2}_{H}}\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\lVert f_{\mu}\rVert_{2,\lambda}^{2}$
$\leq$  $\,9\left(\frac{1}{s_{H}^{4}T^{2}}+\frac{1}{4s^{2}_{H}}\right)\sum_{\mu}\lVert f_{\mu}\rVert_{2,\lambda}^{2}\,.$
The estimate for the part associated with $c_{-}^{\mu}$ can be dealt with similarly. We then conclude
(3.54) $\big{\|}Y^{(0)}_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}\leq\max\left\{\frac{1}{5s^{2}_{H}},\frac{T^{2}}{\pi^{2}}\right\}\lVert F_{t}^{h}\rVert_{2,\lambda\otimes\sigma}^{2}\,,\quad\big{\|}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}\leq 9\left(\frac{4}{s_{H}^{4}T^{2}}+\frac{1}{s^{2}_{H}}\right)\lVert F_{t}^{h}\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
by $\lVert F_{t}^{h}\rVert_{2,\lambda\otimes\sigma}^{2}=\lVert f_{0}\rVert_{2,\lambda}^{2}+\sum_{\mu}\lVert f_{\mu}\rVert_{2,\lambda}^{2}$. In the same manner, we estimate
$\big{\|}\partial_{t}Y^{(0)}_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}\leq\lVert f_{0}\rVert_{2,\lambda}^{2}+2\sum_{\mu}(|c_{+}^{\mu}|^{2}+|c_{-}^{\mu}|^{2})\lVert\psi_{0,\mu}^{\prime}(t)\rVert_{2,\lambda}^{2}\,,$
and
$\big{\|}\partial_{t}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}\leq 2\mu^{2}\sum_{\mu}(|c_{+}^{\mu}|^{2}+|c_{-}^{\mu}|^{2})\lVert\psi_{1,\mu}^{\prime}(t)\rVert_{2,\lambda}^{2}\,.$
By using Eqs. 3.47 and 3.48, as well as Eq. 3.35 with Eq. 3.36, we find
$\sum_{\mu}|c_{+}^{\mu}|^{2}\lVert\psi_{0,\mu}^{\prime}(t)\rVert_{2,\lambda}^{2}\leq$ $\sum_{\mu}\frac{13}{2}\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\lVert f_{\mu}\rVert_{2,\lambda}^{2}$
$\leq$ $\sum_{\mu}\left(\frac{13}{2}+\frac{26}{(\mu T)^{2}}\right)\lVert f_{\mu}\rVert_{2,\lambda}^{2}\,.$
and
$\sum_{\mu}|c_{+}^{\mu}|^{2}\mu^{2}\lVert\psi_{1,\mu}^{\prime}(t)\rVert_{2,\lambda}^{2}\leq$ $\sum_{\mu}\frac{12}{T\mu(1-\ell)v_{T}}\frac{v_{T}^{2}}{v_{T}^{2}-\ell^{2}}\lVert f_{\mu}\rVert_{2,\lambda}^{2}$
$\leq$  $\,36\sum_{\mu}\left(1+\frac{5}{(\mu T)^{2}}+\frac{4}{(\mu T)^{4}}\right)\lVert f_{\mu}\rVert_{2,\lambda}^{2}$
for some universal constant $C$, where the last inequality is implied by
$T\mu(1-\ell)v_{T}=\frac{(1-\ell)^{2}(1+\ell)}{2}\geq\frac{(1-\ell)^{2}}{2}=\frac{(\mu T)^{2}}{2}\,,\quad\text{as}\quad\mu T\to 0\,,$
and hence
$\frac{1}{T\mu(1-\ell)v_{T}}\leq 3+\frac{3}{(\mu T)^{2}}\,.$
Therefore, similarly as for Eq. 3.54, by $\mu^{2}\geq s^{2}_{H}$, we have
$\big{\|}\partial_{t}Y^{(0)}_{t}\big{\|}^{2}_{2,\lambda\otimes\sigma}\leq 52\left(\frac{1}{2}+\frac{2}{(s_{H}T)^{2}}\right)\lVert F_{t}^{h}\rVert_{2,\lambda\otimes\sigma}^{2}\,,$
and
$\big{\|}\partial_{t}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}^{2}\leq 144\left(1+\frac{5}{(s_{H}T)^{2}}+\frac{4}{(s_{H}T)^{4}}\right)\lVert F_{t}^{h}\rVert_{2,\lambda\otimes\sigma}^{2}\,.$
The proof is completed by Eqs. 3.30 and 3.31. ∎
With the help of the above lemmas, we are ready to give the proof of Theorem 3.2.
Proof of Theorem 3.2.
We first note, by triangle inequality,
(3.55) $\lVert X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}\leq\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}+\lVert(1-\Pi_{0})X_{t}\rVert_{2,\lambda\otimes\sigma}\,.$
It suffices to estimate $\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}$. Let $(Y^{(0)}_{t},Y^{(1)}_{t})$ be a solution constructed in Lemma 3.7 for the divergence-type equation (3.26) with source $F_{t}=\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}$. Then, we can compute
(3.56) $\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}^{2}$ $=\left\langle\Pi_{0}X_{t},-\partial_{t}Y^{(0)}_{t}+\left(\mathcal{L}_{+0}^{H}\right)^{\star}Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}$
$=\left\langle\partial_{t}\Pi_{0}X_{t},Y^{(0)}_{t}\right\rangle_{2,\lambda\otimes\sigma}+\left\langle\mathcal{L}_{+0}^{H}\Pi_{0}X_{t},Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}$
$=\left\langle\partial_{t}\Pi_{0}X_{t}-\mathcal{L}^{H}\Pi_{0}X_{t},Y^{(0)}_{t}-Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}\,,$
where the first equality is from
$\left\langle{\bf 1},-\partial_{t}Y^{(0)}_{t}+\left(\mathcal{L}_{+0}^{H}\right)^{\star}Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}=0\,,$
the second equality is by integration by parts for $t$ with $Y_{t}^{(0)}|_{t=0,T}=0$, while the third one uses A and the orthogonality from $Y^{(0)}_{t}\in\mathfrak{H}_{0}$ and $Y^{(1)}_{t}\in\mathfrak{H}_{+}$:
$\left\langle\partial_{t}\Pi_{0}X_{t},Y^{(1)}_{t}\right\rangle_{\sigma,1/2}=\left\langle\mathcal{L}_{+0}^{H}\Pi_{0}X_{t},Y^{(0)}_{t}\right\rangle_{\sigma,1/2}=0\,,\quad\forall\ t\in[0,T]\,.$
We continue to decompose the formula (3.56):
(3.57) $\left\langle(\partial_{t}-\mathcal{L}^{H})\Pi_{0}X_{t},Y^{(0)}_{t}-Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}$
$=$ $\left\langle(\partial_{t}-\mathcal{L}^{H})X_{t},Y^{(0)}_{t}-Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}-\left\langle(\partial_{t}-\mathcal{L}^{H})(1-\Pi_{0})X_{t},Y^{(0)}_{t}-Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma},$
where the second term can be further reformulated as follows:
(3.58) $\Big{\langle}(\partial_{t}-\mathcal{L}^{H})(1-\Pi_{0})X_{t},$ $Y^{(0)}_{t}-Y^{(1)}_{t}\Big{\rangle}_{2,\lambda\otimes\sigma}$
$=\left\langle(1-\Pi_{0})X_{t},\partial_{t}Y^{(1)}_{t}+\mathcal{L}_{+0}^{H}Y^{(0)}_{t}-\mathcal{L}^{H}Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}\,.$
We next estimate $\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\rVert_{2,\lambda\otimes\sigma}$ with the computations in Eqs. 3.56, 3.57 and 3.58. First, thanks to the dissipativity of $\mathcal{L}^{D}$, the operator $\beta-\mathcal{L}^{D}$ is invertible on $L^{2}_{\lambda\otimes\sigma}([0,T];\mathcal{B}(\mathcal{H}))$ for any $\beta>0$, which, along with the estimate (3.27), implies
(3.59) $\left|\left\langle\big{(}\partial_{t}-\mathcal{L}^{H}\big{)}X_{t},Y^{(0)}_{t}-Y^{(1)}_{t}\right\rangle_{2,\lambda\otimes\sigma}\right|$
$\leq$ $\left\lVert\big{(}\beta-\mathcal{L}^{D}\big{)}^{-1/2}\big{(}\partial_{t}-\mathcal{L}^{H}\big{)}X_{t}\right\rVert_{2,\lambda\otimes\sigma}\left\lVert\big{(}\beta-\mathcal{L}^{D}\big{)}^{1/2}\left(Y^{(0)}_{t}-Y^{(1)}_{t}\right)\right\rVert_{2,\lambda\otimes\sigma}$
$\leq$ $\,C_{2,T,\beta}\left\lVert\big{(}\beta-\mathcal{L}^{D}\big{)}^{-1/2}\big{(}\partial_{t}-\mathcal{L}^{H}\big{)}X_{t}\right\rVert_{2,\lambda\otimes\sigma}\left\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\right\rVert_{2,\lambda\otimes\sigma}\,,$
where
$C_{2,T,\beta}=\sqrt{2}\left\lVert(\beta-\mathcal{L}^{D})^{1/2}\right\rVert_{2\to 2}\left(\max\left\{\frac{6}{s_{H}^{2}T}+\frac{3}{s_{H}},\frac{T}{\pi}\right\}+\max\left\{\frac{1}{s_{H}},\frac{T}{\pi}\right\}\right).$
Then, again by estimates in Lemma 3.7, we have
(3.60) $\big{\|}\partial_{t}Y^{(1)}_{t}+\mathcal{L}_{+0}^{H}Y^{(0)}_{t}-\mathcal{L}^{H}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}$
$\leq$ $\big{\|}\partial_{t}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}+\big{\|}\mathcal{L}_{+0}^{H}Y^{(0)}_{t}\big{\|}_{2,\lambda\otimes\sigma}+\big{\|}\mathcal{L}^{H}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}$
$\leq$ $\widetilde{C}_{1,T}\left\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\right\rVert_{2,\lambda\otimes\sigma}\,,$
with
$\widetilde{C}_{1,T}=\left(\sqrt{2}+13+\frac{12\sqrt{5}}{s_{H}T}+\frac{24}{(s_{H}T)^{2}}+\sqrt{2}\max\left\{\frac{6}{s_{H}^{2}T}+\frac{3}{s_{H}},\frac{T}{\pi}\right\}\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}\right)\,,$
where we also used, by similar estimates as Eqs. 3.25, 3.50 and 3.52,
$\big{\|}\mathcal{L}_{+0}^{H}Y^{(0)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\leq\sqrt{2}\left\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\right\rVert_{2,\lambda\otimes\sigma}\,.$
The proof is completed by the following estimate, along with Eqs. 3.55, 3.59 and 3.60,
$\left\lVert\Pi_{0}X_{t}-\langle X_{t}\rangle_{\lambda\otimes\sigma}\right\rVert_{2,\lambda\otimes\sigma}^{2}\leq\big{|}\big{\langle}(\partial_{t}-\mathcal{L}^{H})X_{t},Y^{(0)}_{t}-Y^{(1)}_{t}\big{\rangle}_{2,\lambda\otimes\sigma}\big{|}\\ +\left\lVert(1-\Pi_{0})X_{t}\right\rVert_{2,\lambda\otimes\sigma}\big{\|}\partial_{t}Y^{(1)}_{t}+\mathcal{L}_{+0}^{H}Y^{(0)}_{t}-\mathcal{L}^{H}Y^{(1)}_{t}\big{\|}_{2,\lambda\otimes\sigma}\,.\mbox{\qed}$
We end this section with the proof of Theorem 3.1.
Proof of Theorem 3.1.
Note that if $\langle X_{0}\rangle_{\sigma}=0$, then for any $t>0$, $\langle\mathcal{P}_{t}X_{0}\rangle_{\sigma}=\langle X_{0}\rangle_{\sigma}=0$. Without loss of generality, we assume $\langle X_{0}\rangle_{\sigma}=0$. Then, for any $t,T>0$, we compute
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{1}{T}\int_{t}^{t+T}\lVert X_{\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau$ $=\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{d}t}\lVert X_{t+\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\lambda(\tau)$
(3.61) $=\int_{0}^{T}\langle X_{t+\tau},(\mathcal{L}+\mathcal{L}^{\star})X_{t+\tau}\rangle_{\sigma,1/2}\,\mathrm{d}\lambda(\tau)$
$=2\int_{0}^{T}\langle X_{t+\tau},\mathcal{L}^{D}X_{t+\tau}\rangle_{\sigma,1/2}\,\mathrm{d}\lambda(\tau)=-2\int_{0}^{T}\mathcal{E}_{\mathcal{L}^{D}}(X_{t+\tau},X_{t+\tau})\,\mathrm{d}\lambda(\tau)\,,$
due to
$\mathcal{L}+\mathcal{L}^{\star}=2\mathcal{L}^{D}\,.$
Recalling the coercivity (3.3), we have, for any $s>0$,
$\mathcal{E}_{\mathcal{L}^{D}}(X_{s},X_{s})\geq\lambda_{D}\lVert(1-\Pi_{0})X_{s}\rVert_{\sigma,1/2}^{2}\,.$
On the other hand, by the Lindblad equation,
$\mathcal{L}^{D}X_{s}=(\partial_{s}-\mathcal{L}^{H})X_{s}\,,$
which implies, for any $\beta>0$,
$\mathcal{E}_{\mathcal{L}^{D}}(X_{s},X_{s})$ $=\left\langle\mathcal{L}^{D}X_{s},\left(-\mathcal{L}^{D}\right)^{-1}\mathcal{L}^{D}X_{s}\right\rangle_{\sigma,1/2}$
$\geq\left\lVert\left(\beta-\mathcal{L}^{D}\right)^{-1/2}\mathcal{L}^{D}X_{s}\right\rVert_{\sigma,1/2}^{2}=\left\lVert\left(\beta-\mathcal{L}^{D}\right)^{-1/2}\left(\partial_{s}-\mathcal{L}^{H}\right)X_{s}\right\rVert_{\sigma,1/2}^{2}\,.$
Here $\left(-\mathcal{L}^{D}\right)^{-1}$ is well-defined as the pseudo-inverse of $\mathcal{L}^{D}$. It follows from (3.3) that for any $\eta\in(0,1)$,
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{1}{T}\int_{t}^{t+T}\lVert X_{\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau$
$\leq$ $-2\int_{0}^{T}\eta\lambda_{D}\left\lVert(1-\Pi_{0})X_{t+\tau}\right\rVert_{\sigma,1/2}^{2}+(1-\eta)\left\lVert\left(\beta-\mathcal{L}^{D}\right)^{-1/2}\left(\partial_{\tau}-\mathcal{L}^{H}\right)X_{t+\tau}\right\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\lambda(\tau)\,.$
We take $\eta=\frac{C_{1,T}^{2}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta}^{2}}$ with constants $C_{1,T}$ and $C_{2,T,\beta}$ given in Theorem 3.2, and find
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{1}{T}\int_{t}^{t+T}\lVert X_{\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau\leq$ $-\frac{2\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta}^{2}}\int_{0}^{T}C_{1,T}^{2}\left\lVert(1-\Pi_{0})X_{t+\tau}\right\rVert_{\sigma,1/2}^{2}$
(3.62) $\qquad\qquad\qquad\qquad\quad+C_{2,T,\beta}^{2}\left\lVert\left(\beta-\mathcal{L}^{D}\right)^{-1/2}\left(\partial_{\tau}-\mathcal{L}^{H}\right)X_{t+\tau}\right\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\lambda(\tau)$
$\leq$ $-\frac{\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta}^{2}}\int_{0}^{T}\lVert X_{t+\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\lambda(\tau)\,.$
Then, by Section 3.3 above, a direct application of Grönwall’s inequality gives the exponential decay of the time average of $L^{2}$-distance:
$\frac{1}{T}\int_{t}^{t+T}\lVert X_{\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau\leq e^{-\nu t}\frac{1}{T}\int_{0}^{T}\lVert X_{\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau\,,$
with
(3.63) $\nu=\frac{\lambda_{D}}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta}^{2}}\,.$
The decay estimate in $L^{2}$-distance can be readily implied by the above time-average one:
(3.64) $\lVert X_{t+T}\rVert_{\sigma,1/2}^{2}\leq\frac{1}{T}\int_{t}^{t+T}\lVert X_{\tau}\rVert_{\sigma,1/2}^{2}\,\mathrm{d}\tau\leq e^{-\nu t}\lVert X_{0}\rVert_{\sigma,1/2}^{2}\,,$
since $\lVert X_{t}\rVert_{\sigma,1/2}$ is decreasing in $t$ by the contractivity of QMS. The estimate (3.64) gives
$\lVert X_{t}\rVert_{\sigma,1/2}^{2}\leq e^{-\nu(t-T)}\lVert X_{0}\rVert_{\sigma,1/2}^{2}\,,\quad\forall\,t\geq T\,,$
which in fact holds for any $t>0$, by $\lVert X_{t}\rVert_{\sigma,1/2}\leq\lVert X_{0}\rVert_{\sigma,1/2}$ and $e^{-\nu(t-T)}\geq 1$ for $t\in[0,T]$.
Noting that the convergence rate estimate (3.63)-(3.64) holds for any $\beta>0$, and it is monotonically decreasing in $\beta$, we can take $\beta\to 0$ to maximize the lower bound. Finally, for bounding the prefactor $C_{T}=e^{\nu T}$, it suffices to consider $\nu T$, which can be easily estimated as follows:
$\nu T\leq\frac{\lambda_{D}T}{C_{1,T}^{2}+\lambda_{D}C_{2,T,\beta}^{2}}$ $\leq\frac{\lambda_{D}T\pi^{2}}{(\sqrt{2}+14)^{2}+2\lVert(\mathcal{L}^{D})^{1/2}\rVert_{2\to 2}^{2}\lambda_{D}T^{2}}$
$=\mathcal{O}\left(\frac{\lambda_{D}T}{1+(\lambda_{D}T)^{2}}\right)=\mathcal{O}(1)\,,$
by using Eqs. 3.13 and 3.14 and
$\lVert(\beta-\mathcal{L}^{D})^{1/2}\rVert_{2\to 2}\geq\lVert(\mathcal{L}^{D})^{1/2}\rVert_{2\to 2}=\sup_{X\in\mathfrak{H}}\frac{\lVert(\mathcal{L}^{D})^{1/2}X\rVert_{\sigma,1/2}}{\lVert X\rVert_{\sigma,1/2}}\geq\sqrt{\lambda_{D}}\,.$
The proof is complete. ∎
4. Spectral gap under large coherent limit
Let $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ be a primitive Lindbladian given in (2.5). We have shown (in Corollary 3.4) that if $\mathcal{L}^{D}$ and $\mathcal{L}^{H}$ satisfy A, the lower bound (3.11) of the convergence rate for $\mathcal{L}_{\alpha}$ is increasing in $\alpha>0$. In this section, we consider the asymptotic behavior of the spectral gap $\lambda(\mathcal{L}_{\alpha})$ as $\alpha\to+\infty$, which provides sharper convergence rate estimates. As a corollary, we see that adding a coherent term can accelerate the convergence of Lindblad dynamics in general.
We first reformulate the spectral gap (2.15) as follows:
(4.1) $\lambda(\alpha):=\inf\{\lambda>0\,;\ \exists\mu\in\mathbb{R}\ \,\text{such that}\ -\lambda+i\mu\in{\rm Spec}(\mathcal{L}_{\alpha}|_{\mathfrak{H}})\}\,.$
Suppose that the Hamiltonian $H$ has eigenvalues $\{\lambda_{i}\}$. Then the eigenvalues of the commutator $[H,\cdot]$ are given by the so-called Bohr frequencies:
(4.2) $B_{H}:=\left\{\nu=\lambda_{i}-\lambda_{j}\,;\ \lambda_{i},\lambda_{j}\in{\rm Spec}(H)\right\}\,.$
For each $\nu\in B_{H}$, we denote the associated eigenspace by
$\mathcal{B}_{\nu}:=\{X\in\mathfrak{H}\,;\ [H,X]=\nu X\}\,.$
For ease of exposition, we also introduce the minimal energy of $-\mathcal{L}^{D}$ restricted on $\mathcal{B}_{\nu}$:
(4.3) $\lambda_{\nu}=\inf\{\mathcal{E}_{\mathcal{L}^{D}}(X,X)\,;\ \lVert X\rVert_{\sigma,1/2}=1\,,\ X\in\mathcal{B}_{\nu}\}\,,\quad\nu\in B_{H}\,,$
and the corresponding eigenspace:
(4.4) $\mathcal{M}_{\nu}=\left\{X\in\mathcal{B}_{\nu}\,;\ \mathcal{E}_{\mathcal{L}^{D}}(X,X)=\lambda_{\nu}\lVert X\rVert_{\sigma,1/2}^{2}\right\}\,.$
The main result of this section is Theorem 4.1 below, which is analogous of [franke2010behavior] for the spectral gap of the elliptic equation with growing drift.
Theorem 4.1.
Let $\mathcal{L}_{\alpha}$ be given as in (2.5) and $\lambda(\alpha)$ be the spectral gap given in (4.1). It holds that as $\alpha\to\infty$, the spectral gap $\lambda(\alpha)$ converges to a finite value:
$\lim_{\alpha\to\infty}\lambda(\alpha)=\inf_{\nu\in B_{H}}\lambda_{\nu}\,,$
where $\lambda_{\nu}$ is defined in (4.3).
Remark 9.
We emphasize that we do not assume A for Theorem 4.1, and that by the primitivity of the Lindbladian, each $\lambda_{\nu}$ is positive and hence $\lim_{\alpha\to\infty}\lambda(\alpha)>0$. Indeed, $\lambda_{\nu}>0$ for all Bohr frequencies means that there is no eigenvector of $\mathcal{L}^{H}$ lying in $\ker(\mathcal{L}^{D})$, which is equivalent to the following condition [achleitner2017multi]*Proposition 2.6:
(4.5) $\sum_{j=0}^{J}-(i\mathcal{L}^{H})^{j}\mathcal{L}^{D}(i\mathcal{L}^{H})^{j}\geq K\,{\rm id}\,,\quad\text{on}\ \mathfrak{H}\,,$
for some integer $J\geq 0$ and $K>0$. It can be viewed as a quantum analog of the finite rank Hörmander condition for hypoelliptic degenerate diffusion processes [hormander1967hypoelliptic]. The smallest $J$ such that (4.5) holds is called the hypocoercivity index of $\mathcal{L}_{\alpha}$, noting that $J=0$ gives $-\mathcal{L}^{D}>0$ on $\mathfrak{H}$, that is, $\mathcal{L}_{\alpha}$ is coercive; see Remark 4. Thanks to [achleitner2017multi]*Lemma 2.8, the condition (4.5) holds if and only all the eigenvalues $\lambda$ of $\mathcal{L}_{\alpha}|_{\mathfrak{H}}$ have negative real parts: $\mathfrak{Re}(\lambda)<0$, which is guaranteed by the primitivity (2.6); see also [wolf2012quantum]*Theorem 6.7.
Corollary 4.2 (Accelerating Lindblad dynamics).
Let $\mathcal{L}^{D}$ be a primitive Lindbladian satisfying $\sigma$-KMS DBC, and define $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ with $[H,\sigma]=0$. Then, the spectral gap increases:
(4.6) $\lambda(\alpha)\geq\lambda(\mathcal{L}^{D})\,.$
Suppose that $\mathcal{B}_{D}$ is the eigenspace of $-\mathcal{L}^{D}$ associated with the gap $\lambda(\mathcal{L}^{D})$. The inequality (4.6) strictly holds if and only if $\mathcal{B}_{D}\bigcap\mathcal{B}_{\nu}=\{0\}$ for any $\nu\in B_{H}$. In this case, we still have the strict increase of the spectral gap in the limit:
$\lim_{\alpha\to\infty}\lambda(\alpha)>\lambda(\mathcal{L}^{D})\,.$
Proof.
The inequality (4.6) simply restates Eq. 2.20 in Remark 4. Now, suppose $\lambda(\alpha)=\lambda(\mathcal{L}^{D})$. Let $X_{\alpha}$ be an eigenvector of $\mathcal{L}_{\alpha}$ satisfying $\mathcal{L}_{\alpha}X_{\alpha}=\lambda X_{\alpha}$ with $\mathfrak{Re}(\lambda)=-\lambda(\alpha)=-\lambda(\mathcal{L}^{D})$. The argument for Eq. 2.20, along with Min-Max theorem, gives $\mathcal{L}^{D}(X_{\alpha})=-\lambda(\mathcal{L}^{D})X_{\alpha}$, which implies that $X_{\alpha}$ is also an eigenvector of $\mathcal{L}^{H}$. This shows the only if direction, while the if direction is trivial. The proof is completed by Theorem 4.1. ∎
The proof of Theorem 4.1 relies on the following two propositions.
Proposition 4.3.
Suppose that $\limsup_{\alpha\to\infty}\lambda(\alpha)<\infty$. There holds
$\liminf_{\alpha\to\infty}\lambda(\alpha)\geq\inf_{\nu\in B_{H}}\lambda_{\nu}\,.$
Proposition 4.4.
Given $\alpha>0$ and a Bohr frequency $\nu\in B_{H}$, let $\lambda_{\nu}$ be given in (4.3) and denote by $B_{\eta}(-\lambda_{\nu}+i\alpha\nu)$ the $\eta$-neighborhood of $-\lambda_{\nu}+i\alpha\nu$ in the complex plane:
$B_{\eta}(-\lambda_{\nu}+i\alpha\nu):=\{z\in\mathbb{C}\,;\ |z-(-\lambda_{\nu}+i\alpha\nu)|\leq\eta\}\,.$
Then, for any $\eta>0$, there exists $\alpha_{0}$ large enough such that for all $\alpha\geq\alpha_{0}$,
${\rm Spec}(\mathcal{L}_{\alpha}|\mathfrak{H})\bigcap B_{\eta}(-\lambda_{\nu}+i\alpha\nu)\neq\emptyset.$
Proof of Theorem 4.1.
It suffices to note from Proposition 4.4 that for any $\nu\in B_{H}$, there holds
$\limsup_{\alpha\to\infty}\lambda(\alpha)\leq\lambda_{\nu}\,,$
where $\lambda_{\nu}$ is defined in (4.3). Then the proof is readily completed by Proposition 4.3. ∎
The rest of this section is devoted to the proofs of Propositions 4.3 and 4.4. We start with the easier lower bound of the limiting spectral gap (Proposition 4.3).
Proof of Proposition 4.3.
Suppose that $-\lambda(\alpha)+i\mu_{\alpha}\in\mathbb{C}$ is the eigenvalue of $\mathcal{L}_{\alpha}$ with $\lambda(\alpha)$ being the spectral gap (4.1), and $X^{\alpha}=X_{1}^{\alpha}+iX_{2}^{\alpha}\in\mathfrak{H}$ with $\lVert X^{\alpha}\rVert_{\sigma,1/2}=1$ is the associated eigenvector:
(4.7) $\mathcal{L}_{\alpha}(X_{1}^{\alpha}+iX_{2}^{\alpha})=(-\lambda(\alpha)+i\mu_{\alpha})X^{\alpha}\,,$
where $X_{1}^{\alpha},X_{2}^{\alpha}\in\mathcal{B}(\mathcal{H})$ are real-valued. It follows that
$-\langle X^{\alpha},\mathcal{L}_{\alpha}X^{\alpha}\rangle_{\sigma,1/2}=\lambda(\alpha)-i\mu_{\alpha}\,.$
Since $\mathcal{L}^{H}$ is anti-Hermitian, we readily have
(4.8) $\mathcal{E}_{\mathcal{L}^{D}}(X^{\alpha},X^{\alpha})=-\mathfrak{Re}\,\langle X^{\alpha},\mathcal{L}_{\alpha}X^{\alpha}\rangle=\lambda(\alpha)\,.$
In addition, by the boundedness of $\lVert X^{\alpha}\rVert_{\sigma,1/2}$, there is a convergent subsequence of $\{X^{\alpha}\}$ (without loss of generality, still denoted by $X^{\alpha}$) such that
$X_{*}:=\lim_{\alpha\to\infty}X^{\alpha}\in\mathfrak{H}\,.$
Note from Eq. 4.7 that for any $Y\in\mathfrak{H}$,
$\mathcal{E}_{\mathcal{L}_{\alpha}}(Y,X^{\alpha})=(\lambda(\alpha)-i\mu_{\alpha})\langle Y,X^{\alpha}\rangle_{\sigma,1/2}\,,$
which gives
$-i\langle Y,[H,X^{\alpha}]\rangle_{\sigma,1/2}+\frac{1}{\alpha}\mathcal{E}_{\mathcal{L}^{D}}(Y,X^{\alpha})=\frac{1}{\alpha}(\lambda(\alpha)-i\mu_{\alpha})\langle Y,X^{\alpha}\rangle_{\sigma,1/2}\,.$
Then, letting $\alpha\to\infty$ and recalling $\lim\sup_{\alpha\to\infty}\lambda(\alpha)<\infty$, we find
$i\langle Y,[H,X_{*}]\rangle_{\sigma,1/2}=\left(\lim_{\alpha\to\infty}\frac{i\mu_{\alpha}}{\alpha}\right)\langle Y,X_{*}\rangle_{\sigma,1/2}\,,$
which means that $\nu_{*}:=\lim_{\alpha\to\infty}\frac{\mu_{\alpha}}{\alpha}$ exists and is a Bohr frequency with $X_{*}\in\mathcal{B}_{\nu_{*}}$:
$\mathcal{L}^{H}X_{*}=i\nu_{*}X_{*}\,.$
The proof is completed by (4.8):
$\liminf_{\alpha\to\infty}\lambda(\alpha)\geq\inf\{\mathcal{E}_{\mathcal{L}^{D}}(X,X)\,;\ \lVert X\rVert_{\sigma,1/2}=1\,,\ X\in\mathcal{B}_{\nu}\ \text{for some $\nu$}\}\,.\qed$
We next prove Proposition 4.4 on the stability of the spectrum of $\mathcal{L}_{\alpha}$.
Proof of Proposition 4.4.
We proceed by a contradiction argument. Suppose that there exists $\eta>0$ and $\alpha_{n}\to\infty$ such that any element $z$ in the set
$B_{\eta}(-\lambda_{\nu}+i\alpha_{n}\nu)=\{-\lambda+i\mu\,;\ \lambda=\lambda_{\nu}+\varepsilon\in\mathbb{R}\,,\ \mu=\alpha_{n}\nu+\delta\in\mathbb{R}\,,\quad\varepsilon^{2}+\delta^{2}\leq\eta^{2}\}$
is not an eigenvalue of $\mathcal{L}_{\alpha}|_{\mathfrak{H}}$. We fix $Y\in\mathcal{M}_{\nu}$ (4.1) with normalization $\lVert Y\rVert_{\sigma,1/2}=1$. Then, for any $z=-\lambda+i\mu\in B_{\eta}(-\lambda_{\nu}+i\alpha_{n}\nu)$, there exists $X^{\alpha_{n}}_{z}\in\mathfrak{H}$ satisfying
(4.9) $(z-\mathcal{L}_{\alpha_{n}})X^{\alpha_{n}}_{z}=Y\,,$
since $z-\mathcal{L}_{\alpha_{n}}$ is invertible. Note that $z\in\mathbb{C}$ is determined by $(\varepsilon,\delta)$ and $-\lambda_{\nu}+i\alpha_{n}\nu$.
We first show that $\{X_{z}^{\alpha_{n}}\}$ is an uniformly bounded sequence: for small enough $\eta$,
(4.10) $\limsup_{n\to\infty}\sup_{\varepsilon^{2}+\delta^{2}=\eta^{2}}\lVert X^{\alpha_{n}}_{z}\rVert_{\sigma,1/2}<\infty\,.$
If it is not true, there exists a sequence of real $\varepsilon_{n},\delta_{n}$ such that $\varepsilon_{n}^{2}+\delta_{n}^{2}=\eta^{2}$, $z_{n}=-(\lambda_{\nu}+\varepsilon_{n})+i(\alpha_{n}\nu+\delta_{n})$, and when $n\to\infty$,
$K_{n}:=\lVert X^{\alpha_{n}}_{z_{n}}\rVert_{\sigma,1/2}\to\infty\,.$
We define $\widetilde{X}_{n}=X^{\alpha_{n}}_{z_{n}}/K_{n}$, which is bounded and satisfies, by (4.9), for any $Z\in\mathfrak{H}$,
(4.11) $z_{n}\langle Z,\widetilde{X}_{n}\rangle_{\sigma,1/2}-i\alpha_{n}\langle Z,[H,\widetilde{X}_{n}]\rangle_{\sigma,1/2}+\mathcal{E}_{\mathcal{L}^{D}}(Z,\widetilde{X}_{n})=\frac{1}{K_{n}}\langle Z,Y\rangle_{\sigma,1/2}\,.$
By the boundedness of $\widetilde{X}_{n}$, without loss of generality, we can assume that as $n\to\infty$, for some $\widetilde{X}_{*}\in\mathfrak{H}$ and real $\varepsilon_{*},\delta_{*}$ satisfying $\varepsilon_{*}^{2}+\delta_{*}^{2}=\eta^{2}$, there holds
$\widetilde{X}_{n}\to\widetilde{X}_{*}\,,\quad\varepsilon_{n}\to\varepsilon_{*}\,,\quad\delta_{n}\to\delta_{*}\,.$
Then, from Eq. 4.11, we have
$\frac{z_{n}}{\alpha_{n}}\langle Z,\widetilde{X}_{n}\rangle_{\sigma,1/2}-i\langle Z,[H,\widetilde{X}_{n}]\rangle_{\sigma,1/2}+\frac{1}{\alpha_{n}}\mathcal{E}_{\mathcal{L}^{D}}(Z,\widetilde{X}_{n})=\frac{1}{\alpha_{n}K_{n}}\langle Z,Y\rangle_{\sigma,1/2}\,,$
and let $n\to\infty$ to obtain
$i\nu\langle Z,\widetilde{X}_{*}\rangle_{\sigma,1/2}-i\langle Z,[H,\widetilde{X}_{*}]\rangle_{\sigma,1/2}=0\,,$
that is, $\widetilde{X}_{*}\in\mathcal{B}_{\nu}$. We now consider test operators $Z=\widetilde{X}_{n}$ in (4.11) and take the real part:
$-(\lambda_{\nu}+\varepsilon_{n})\lVert\widetilde{X}_{n}\rVert_{\sigma,1/2}^{2}+\mathcal{E}_{\mathcal{L}^{D}}(\widetilde{X}_{n},\widetilde{X}_{n})=\frac{1}{K_{n}}\mathfrak{Re}\,\langle\widetilde{X}_{n},Y\rangle_{\sigma,1/2}\,,$
which implies, by letting $n\to\infty$,
(4.12) $\mathcal{E}_{\mathcal{L}^{D}}(\widetilde{X}_{*},\widetilde{X}_{*})=(\lambda_{\nu}+\varepsilon_{*})\lVert\widetilde{X}_{*}\rVert_{\sigma,1/2}^{2}\,.$
It follows that $\varepsilon_{*}\geq 0$ by $\widetilde{X}_{*}\in\mathcal{B}_{\nu}$ and the definition (4.3) of $\lambda_{\nu}$. On the other hand, we take test operators $Z\in\mathcal{M}_{\nu}$ (4.4) in (4.11). A similar calculation gives
$(\lambda_{\nu}-i\alpha_{n}\nu+z_{n})\langle Z,\widetilde{X}_{n}\rangle_{\sigma,1/2}=\frac{1}{K_{n}}\langle Z,Y\rangle_{\sigma,1/2}\,,$
which further yields $(-\varepsilon_{*}+i\delta_{*})\langle Z,\widetilde{X}_{*}\rangle_{\sigma,1/2}=0$ by $n\to\infty$, and hence
$\widetilde{X}_{*}\perp\mathcal{M}_{\nu}\,.$
By the Min-Max theorem for the eigenvalue problem, it holds that $\mathcal{E}_{\mathcal{L}^{D}}(\widetilde{X}_{*},\widetilde{X}_{*})/\lVert\widetilde{X}_{*}\rVert_{\sigma,1/2}^{2}$ is strictly greater than the second smallest eigenvalue of $-\mathcal{L}^{D}|_{\mathcal{B}_{\nu}}$. This contradicts with (4.12) as $\varepsilon_{*}$ could be arbitrarily small. We have proved the claim (4.10).
Next, we fix a small enough $\eta$. By the boundedness of $X^{\alpha_{n}}_{z_{n}}$ proved in (4.10), up to a subsequence, we have $X^{\alpha_{n}}_{z_{n}}\to X_{*}$ for some $X_{*}\in\mathfrak{H}$ as $n\to\infty$, where $X^{\alpha_{n}}_{z_{n}}$ solves the equation (4.9) and
$z_{n}=-(\lambda_{\nu}+\varepsilon)+i(\alpha_{n}\nu+\delta)$
with $\varepsilon^{2}+\delta^{2}=\eta^{2}$. Similarly to the argument for (4.10), we consider the equation (4.9) with test operators $Z=Y\in\mathcal{M}_{\nu}$ and find
$z_{n}\langle Y,X^{\alpha_{n}}_{z_{n}}\rangle_{\sigma,1/2}-i\alpha_{n}\langle Y,[H,X^{\alpha_{n}}_{z_{n}}]\rangle_{\sigma,1/2}+\mathcal{E}_{\mathcal{L}^{D}}(Y,X^{\alpha_{n}}_{z_{n}})=\langle Y,Y\rangle_{\sigma,1/2}\,,$
and by letting $n\to\infty$,
(4.13) $(-\varepsilon+i\delta)\langle Y,X_{*}\rangle_{\sigma,1/2}=1\,.$
The same analysis also shows that for $Z\in\mathcal{M}_{\nu}\perp Y$, there holds
$\langle Z,X_{*}\rangle_{\sigma,1/2}=0\,.$
This means that the projection of $X_{*}$ on $\mathcal{M}_{\nu}$ is given by $-\frac{\varepsilon+i\delta}{\varepsilon^{2}+\delta^{2}}Y$. We define
(4.14) $\widehat{X}_{*}=X_{*}+\frac{\varepsilon+i\delta}{\varepsilon^{2}+\delta^{2}}Y\,.$
Moreover, we divide Eq. 4.9 with $z_{n}$ by $\alpha_{n}$ and obtain $X_{*}\in\mathcal{B}_{\nu}$ by letting $n\to\infty$. Then, taking test operators $Z=X^{\alpha_{n}}_{z_{n}}$ for the equation (4.9), we have
$-(\lambda_{\nu}+\varepsilon)\langle X_{*},X_{*}\rangle_{\sigma,1/2}+\mathcal{E}_{\mathcal{L}^{D}}(X_{*},X_{*})=\mathfrak{Re}\,\langle X_{*},Y\rangle_{\sigma,1/2}=-\frac{\varepsilon}{\varepsilon^{2}+\delta^{2}}\,,$
by (4.13). Plugging (4.14) into the above formula gives, by orthogonality between $\widehat{X}_{*}$ and $Y$,
$-(\lambda_{\nu}+\varepsilon)\left(\lVert\widehat{X}_{*}\rVert_{\sigma,1/2}^{2}+\frac{1}{\varepsilon^{2}+\delta^{2}}\right)+\mathcal{E}_{\mathcal{L}^{D}}(\widehat{X}_{*},\widehat{X}_{*})+\frac{\lambda_{\nu}}{\varepsilon^{2}+\delta^{2}}=-\frac{\varepsilon}{\varepsilon^{2}+\delta^{2}}\,,$
which implies
$\mathcal{E}_{\mathcal{L}^{D}}(\widehat{X}_{*},\widehat{X}_{*})=(\lambda_{\nu}+\varepsilon)\lVert\widehat{X}_{*}\rVert_{\sigma,1/2}^{2}\,.$
Again, thanks to $\widehat{X}_{*}\perp\mathcal{M}_{\nu}$ and that $\varepsilon$ could be arbitrarily small, it holds that
$\widehat{X}_{*}=0\quad\text{and}\quad X_{*}=-\frac{\varepsilon+i\delta}{\varepsilon^{2}+\delta^{2}}Y\,.$
We are now ready to complete the proof. By assumption, for any small $\eta$, there is no spectrum of $\mathcal{L}_{\alpha}|_{\mathfrak{H}}$ in $B_{\eta}(-\lambda_{v}+i\alpha_{n}\nu)$. It follows that
$\int_{\Gamma_{c}}(z-\mathcal{L}_{\alpha_{n}}|_{\mathfrak{H}})^{-1}Y\,\mathrm{d}z=0\,,$
where $\Gamma_{c}=\{z\in\mathbb{C}\,;\ |z-(-\lambda_{v}+i\alpha_{n}\nu)|=\eta\}$. Then, we have
$\lVert 2\pi Y\rVert_{\sigma,1/2}^{2}$ $=\Big{\|}\int_{\Gamma_{c}}(z-\mathcal{L}_{\alpha_{n}}|_{\mathfrak{H}})^{-1}Y\,\mathrm{d}z-\int_{\Gamma_{c}}\frac{1}{z-(-\lambda_{v}+i\alpha_{n}\nu)}\,\mathrm{d}zY\Big{\|}_{\sigma,1/2}^{2}$
(4.15) $\leq\int_{\Gamma_{c}}\Big{\|}(z-\mathcal{L}_{\alpha_{n}}|_{\mathfrak{H}})^{-1}Y-\frac{1}{z-(-\lambda_{v}+i\alpha_{n}\nu)}Y\Big{\|}_{\sigma,1/2}^{2}\,\mathrm{d}z$
$\leq\int_{\Gamma_{c}}\Big{\|}X^{\alpha_{n}}_{z}+\frac{\varepsilon+i\delta}{\varepsilon^{2}+\delta^{2}}Y\Big{\|}_{\sigma,1/2}^{2}\,\mathrm{d}z\,.$
Recall that we have proved the boundedness of $X^{\alpha_{n}}_{z}$ with fixed $\varepsilon$ and $\delta$, and it converges to $-\frac{\varepsilon+i\delta}{\varepsilon^{2}+\delta^{2}}Y$ as $n\to\infty$. Combining this with (4), we have $Y=0$, which contradicts with $\lVert Y\rVert_{\sigma,1/2}=1$. ∎
5. Applications and examples
In this section, we apply the results in Section 3 to Davies generators with detailed balance broken by a coherent term, namely, the Lindbladian $\mathcal{L}_{\alpha}$ in (2.5) with $\mathcal{L}^{D}$ satisfying $\sigma$-GNS DBC (noting that up to a coherent term, the Davies generator is exactly the class of $\sigma$-GNS detailed balanced Lindbladian [kossakowski1977quantum]; see also [ding2024efficient]*Section 2.1). For the reader’s convenience, we first recall the canonical form for the GNS detailed balanced QMS [alicki1976detailed].
Lemma 5.1.
For a given full-rank state $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, a Lindbladian $\mathcal{L}$ satisfying $\sigma$-GNS DBC is of the following form:
(5.1) $\mathcal{L}(X)=\sum_{j\in\mathcal{J}}\left(e^{-\omega_{j}/2}L_{j}^{\dagger}[X,L_{j}]+e^{\omega_{j}/2}[L_{j},X]L^{\dagger}_{j}\right)\,,$
with $\omega_{j}\in\mathbb{R}$ and $|\mathcal{J}|\leq N^{2}-1$, where jumps $L_{j}\in\mathcal{B}(\mathcal{H})$ satisfy
(5.2) $\Delta_{\sigma}(L_{j})=e^{-\omega_{j}}L_{j}\,,\quad\operatorname{tr}(L^{\dagger}_{j}L_{k})=c_{j}\delta_{j,k}\,,\quad\operatorname{tr}(L_{j})=0\,,$
for some normalization constants $c_{j}>0$. Moreover, for each $j$, there exists $j^{\prime}\in\mathcal{J}$ such that
(5.3) $L_{j}^{\dagger}=L_{j^{\prime}}\,,\quad\omega_{j}=-\omega_{j^{\prime}}\,.$
With the help of Lemma 5.1, it is easy to design some primitive hypocoercive Lindblad dynamics with broken GNS detailed balance. Specifically, we start with a non-primitive $\mathcal{L}^{D}$ of the form (5.1) with $\sigma$-GNS DBC. From Lemma 2.2, its kernel is given by $\ker(\mathcal{L}^{D})=\{X\,;\ [L_{j},X]=0\ \text{for all $j\in\mathcal{J}$}\}$. In some cases, it is possible to efficiently characterize $\ker(\mathcal{L}^{D})$, allowing the design of a coherent term $\mathcal{L}^{H}=i[H,\cdot]$ such that $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ is primitive with A. Consequently, the results of Section 3 can be applied to quantify the convergence rate of the Lindblad dynamics $e^{t\mathcal{L}_{\alpha}}$.
5.1. Infinite temperature Gibbs state
Let us first consider the simpler infinite temperature regime, where $\sigma={\bf 1}/N$ is the maximally mixed state. In this case, the KMS inner product $\langle\cdot,\cdot\rangle_{\sigma,1/2}$ reduces to the normalized Hilbert-Schmidt inner product:
(5.4) $\langle X,Y\rangle_{\frac{{\bf 1}}{N}}:=\frac{1}{N}\langle X,Y\rangle\,,$
and $\Delta_{\sigma}$ becomes the identity operator. This implies $\omega_{j}=0$ in (5.1) and that the jumps $L_{j}$ can be taken to be self-adjoint. Thus, the Lindbladian with $\frac{{\bf 1}}{N}$-GNS DBC satisfies $\mathcal{L}=\mathcal{L}^{\dagger}$ and has the following general form: for some self-adjoint operators $L_{j}$,
(5.5) $\mathcal{L}(X):=-\sum_{j\in\mathcal{J}}[L_{j},[L_{j},X]]\,.$
We shall focus on the dissipative Lindbladian $\mathcal{L}^{D}$ with a single jump to illustrate the ideas and leave the general case (5.5) to future work. Let $A$ be a self-adjoint operator on $\mathcal{H}$ with spectral decomposition: $A=\sum_{i}\kappa_{i}P_{i}$, where $P_{i}$ is the spectral projection associated with the eigenvalue $\kappa_{i}$. We define the non-primitive Lindbladian with jump $A$:
(5.6) $\mathcal{L}^{D}(X):=-[A,[A,X]]=-\sum_{i,j}(\kappa_{i}-\kappa_{j})^{2}P_{i}XP_{j}\,,$
which is self-adjoint on $\mathcal{B}(\mathcal{H})$. The eigenvalues of $\mathcal{L}^{D}$ are given by $-(\kappa_{i}-\kappa_{j})^{2}$ with the eigenspaces $P_{i}\mathcal{B}(\mathcal{H})P_{j}$. We hence see the norm and spectral gap of $\mathcal{L}^{D}$:
(5.7) $\lVert\mathcal{L}^{D}\rVert_{2\to 2}=\max_{i,j}|\kappa_{i}-\kappa_{j}|^{2}\,,\quad\lambda_{D}=\min_{i\neq j}|\kappa_{i}-\kappa_{j}|^{2}\,,$
as well as the kernel $\ker(\mathcal{L}^{D})=\{X\,;\ [A,X]=0\}$. In what follows, for simplicity, we assume that each $\kappa_{j}$ is of multiplicity one, and we denote $P_{j}=\lvert j\rangle\langle j\rvert$. Then $\ker(\mathcal{L}^{D})$ is characterized by the diagonal matrices on the basis $\lvert i\rangle\langle j\rvert$:
(5.8) $\ker(\mathcal{L}^{D})=\Big{\{}X\in\mathcal{B}(\mathcal{H})\,;X=\sum_{j}x_{j}\lvert j\rangle\langle j\rvert\,,\ x_{j}\in\mathbb{C}\Big{\}}\,.$
We next characterize the coherent term $\mathcal{L}^{H}=i[H,\cdot]$ such that $\ker(\mathcal{L}^{H})\bigcap\ker(\mathcal{L}^{D})={\rm Span}\{{\bf 1}\}$, which by Lemma 2.3, implies that $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ is primitive. We write $H=\sum_{ij}H_{ij}\lvert i\rangle\langle j\rvert$, and then by Eq. 5.8, $X\in\ker(\mathcal{L}^{H})\bigcap\ker(\mathcal{L}^{D})$ is equivalent to
(5.9) $(x_{i}-x_{j})H_{ij}=0\,.$
The condition $\ker(\mathcal{L}^{H})\bigcap\ker(\mathcal{L}^{D})={\rm Span}\{{\bf 1}\}$ implies that the constant vector $(x_{j})_{j}$ is the only solution to Eq. 5.9. It follows that for any two distinct indices $i_{1},i_{m}$, there exists a sequence of indices $(i_{1},i_{2},\ldots,i_{m})$ such that $H_{i_{k}i_{k+1}}\neq 0$ for $11\leq k\leq m-$ 1, that is, $H_{ij}$ is an irreducible self-adjoint matrix. We proceed to characterize $H$ satisfying A: $\Pi_{0}\mathcal{L}^{H}\Pi_{0}=0$, where $\Pi_{0}$ is the projection to $\ker(\mathcal{L}^{D})$. By our construction, for any diagonal matrices $X=\sum_{j}x_{j}\lvert j\rangle\langle j\rvert$ and $H$,
$\langle X,[H,X]\rangle_{\frac{{\bf 1}}{N}}=0\,,$
i.e., $\Pi_{0}\mathcal{L}^{H}\Pi_{0}=0$ always holds. Note that $\mathfrak{H}_{0}$ in (3.2) in this example is parameterized by
$\mathfrak{H}_{0}=\Big{\{}X\in\mathcal{B}(\mathcal{H})\,;X=\sum_{j}x_{j}\lvert j\rangle\langle j\rvert\,,\ \sum_{j}x_{j}=0\Big{\}},$
and there holds
$\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}\leq\lambda_{\max}(H)-\lambda_{\min}(H)\,.$
To apply Corollary 3.3, it suffices to estimate
(5.10) $s_{H}^{2}=\inf_{X\in\mathfrak{H}_{0}}\frac{\lVert\mathcal{L}^{H}\Pi_{0}X\rVert_{\frac{{\bf 1}}{N}}^{2}}{\lVert X\rVert_{\frac{{\bf 1}}{N}}^{2}}\,,$
where for $X\in\mathfrak{H}_{0}$, $\lVert X\rVert_{\frac{{\bf 1}}{N}}^{2}=\frac{1}{N}\sum_{j}|x_{j}|^{2}$ and
(5.11) $\lVert\mathcal{L}^{H}\Pi_{0}X\rVert_{\frac{{\bf 1}}{N}}^{2}=\langle[H,X],[H,X]\rangle_{\frac{{\bf 1}}{N}}=\frac{1}{N}\sum_{i,j}|x_{j}-x_{i}|^{2}|H_{ij}|^{2}\,.$
Since the values of $|H_{jj}|^{2}$ do not influence the sum (5.11), we define the symmetric matrix $\widehat{H}$ by
(5.12) $\widehat{H}_{ij}:=|H_{ij}|^{2}-\delta_{ij}\sum_{l}|H_{il}|^{2}\,,$
which is non-negative off the diagonal and still irreducible. Moreover, each row of $\widehat{H}$ sums to zero. Thus, $\widehat{H}$ can be viewed as the generator of an ergodic continuous-time Markov chain over the basis $\lvert j\rangle$ and Eq. 5.10 gives
(5.13) $s_{H}^{2}=\inf_{x_{j}\in\mathbb{C};\,\sum_{j}x_{j}=0}\frac{\sum_{i,j}|x_{j}-x_{i}|^{2}\widehat{H}_{ij}}{\sum_{j}|x_{j}|^{2}}=\inf_{x_{j}\in\mathbb{C};\,\sum_{j}x_{j}=0}\frac{-2\big{\langle}x,\widehat{H}x\big{\rangle}}{\langle x,x\rangle}\,,$
where $x=(x_{1},\ldots,x_{N})\in\mathbb{C}^{N}$. It follows that $s_{H}/\sqrt{2}$ is the square root of the spectral gap (smallest nonzero eigenvalue) of $-\widehat{H}$. Summarizing the above discussion and applying Corollary 3.3 with Eq. 3.17, we have the following result.
Proposition 5.2.
Let $\mathcal{L}^{D}$ be given as in (5.6) for a self-adjoint $A=\sum_{j}\kappa_{j}\lvert j\rangle\langle j\rvert$ with eigenvalue $\kappa_{j}$ of multiplicity one. Then, for a Hamiltonian $H=\sum_{ij}H_{ij}\lvert i\rangle\langle j\rvert$, the Lindbladian $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ is primitive if and only if $H_{ij}$ is irreducible. In this case, the exponential decay holds: for $X_{0}\in\mathcal{B}(\mathcal{H})$ and $X_{t}=e^{t\mathcal{L}_{\alpha}}X_{0}$,
$\lVert X_{t}-\langle X_{t}\rangle_{\frac{{\bf 1}}{N}}\rVert_{\frac{{\bf 1}}{N}}^{2}\leq Ce^{-\nu t}\lVert X_{0}-\langle X_{0}\rangle_{\frac{{\bf 1}}{N}}\rVert_{\frac{{\bf 1}}{N}}^{2}\,,\quad\forall\,t\geq 0\,,$
with convergence rate:
(5.14) $\nu$ $=\frac{\lambda(-\widehat{H})\min_{i\neq j}|\kappa_{i}-\kappa_{j}|^{2}}{\bigl{(}28\lambda(-\widehat{H})^{1/2}+5\bigl{(}\lambda_{\max}(H)-\lambda_{\min}(H)\bigr{)}\bigr{)}^{2}+36\alpha^{-2}\min_{i\neq j}|\kappa_{i}-\kappa_{j}|^{2}\max_{i,j}|\kappa_{i}-\kappa_{j}|^{2}}$
and the prefactor $C=\mathcal{O}(1)$, where $\lambda(-\widehat{H})$ denotes the spectral gap of $-\widehat{H}$ in (5.12).
Recalling Remark 8, it is useful to find a lower bound of $\lambda(-\widehat{H})$ to derive a more concrete estimate. For this, we identify $\widehat{H}$ with a weighted directed graph $G$ on vertices $j$ with oriented edges $(i,j)$ of weights $\widehat{H}_{ij}$ connecting $i$ and $j$ if $\widehat{H}_{ij}>0$. Such a graph $G$ uniquely specifies the matrix $\widehat{H}$ as the graph Laplacian. This allows us to estimate the spectral gap of $\widehat{H}$ (and $s_{H}$) by the canonical path argument [fill1991eigenvalue, diaconis1991geometric, sinclair1992improved] utilizing the geometry of the underlying graph $G$. Indeed, following [sinclair1992improved], a canonical path $\gamma_{ij}$ in $G$ is a path from $i$ to $j$, where vertices can be repeated but no edge is overloaded. We denote by $|\gamma_{ij}|$ the length of the path, which is equal to the number of edges included in $\gamma_{ij}$. Note that the infimum in (5.13) can be taken over real $\{x_{j}\}$, since $\widehat{H}$ is Hermitian. Then, by the proof of [sinclair1992improved]*Theorem 5, for any $x\in\mathbb{R}^{N}$ with $\sum_{j}x_{j}=0$ and any collection of canonical paths $\Gamma=\{\gamma_{ij}\}_{i\neq j}$, we have
(5.15) $2\sum_{j}x_{j}^{2}=\sum_{i\neq j}(x_{i}-x_{j})^{2}$ $=\sum_{i\neq j}\Big{(}\sum_{e\in\gamma_{ij}}(x_{e^{+}}-x_{e^{-}})\Big{)}^{2}$
$\leq\sum_{i\neq j}|\gamma_{ij}|\sum_{e\in\gamma_{ij}}(x_{e^{+}}-x_{e^{-}})^{2}$
$=\sum_{e}(x_{e^{+}}-x_{e^{-}})^{2}\widehat{H}_{e^{+}e^{-}}\sum_{e\in\gamma_{ij}}\big{(}\widehat{H}_{e^{+}e^{-}}\big{)}^{-1}|\gamma_{ij}|\,,$
where $e^{\pm}$ denote the end points of an edge $e=(e^{+},e^{-})$. We define the quantity
$K_{\widehat{H},\Gamma}=\max_{e=(e^{+},e^{-})}\big{(}\widehat{H}_{e^{+}e^{-}}\big{)}^{-1}\sum_{e\in\gamma_{ij}}|\gamma_{ij}|\,,$
depending on $\widehat{H}$ and the choice of paths $\Gamma$. It follows from Eqs. 5.13 and 5.15 that
(5.16) $s_{H}=\sqrt{2}\lambda(-\widehat{H})\geq\sqrt{2/K_{\widehat{H},\Gamma}}\,.$
Plugging it into Eq. 5.14 gives a more manageable lower bound of the convergence rate $\nu$.
Remark 10.
It is known [sinclair1992improved] that $K_{\widehat{H},\Gamma}$ measures the “bottleneck” of the graph $G$. If one could choose canonical paths that do not pass any single edge too many times, the quantity $K_{\widehat{H},\Gamma}$ would be small and we could find a tighter lower for $s_{H}$ and $\nu$. See [sinclair1992improved, diaconis1991geometric] for examples.
We should point out that using $K_{\widehat{H},\Gamma}$ is just one of many possible approaches to bound the spectral gap of $\widehat{H}$ and $s_{H}$. For example, by [diaconis1991geometric], one can also estimate $s_{H}$ in terms of
$\tau_{\widehat{H},\Gamma}:=\max_{e}\sum_{e\in\gamma_{ij}}|\gamma_{ij}|_{\widehat{H}}\quad\text{with}\quad|\gamma_{ij}|_{\widehat{H}}:=\sum_{e\in\gamma_{ij}}\big{(}\widehat{H}_{e^{+}e^{-}}\big{)}^{-1}\,.$
It is clear that if $\widehat{H}_{ij}$ is constant (which is the case when $\widehat{H}$ corresponds to a random walk on a graph), then $\tau_{\widehat{H},\Gamma}=K_{\widehat{H},\Gamma}$. However, in general, these two quantities are not comparable. Other techniques for estimating the spectral gap of a classical Markov chain, including Cheeger’s inequality, coupling, and stopping times (see e.g., the monograph [levin2017markov]) can also be adapted to our case.
5.2. Finite temperature Gibbs state
We next consider examples for the finite temperature Gibbs state preparation via the irreversible Lindbladian (2.5). We shall focus on a quantum analog of the weighted graph Laplacian with detailed balance condition broken by a coherent term, inspired by the models in [gao2022complete]*Section 5 and [li2020graph].
Let $G=(V,E)$ be a disconnected undirected simple graph [^2] A graph without loops or multiple edges is called a simple graph. with vertices $V:=\{1,2,\ldots,N\}$ and edge set $E$. Suppose that $G$ has no isolated vertex (i.e., each node $r\in V$ is connected to at least one other node $s\neq r$). For convenience, we identify the node $r$ with the basis vector $\lvert r\rangle$ of $\mathbb{C}^{N}$ with $1$ at $r$th position and 00 elsewhere. Without loss of generality, by reordering the vertices, we assume that the connected components of $G=(V,E)$ are given by
(5.17) $\{G_{i}=(V_{i},E_{i})\}_{i=1}^{m}\,,\quad 1\leq m\leq[N/2]\,,$
with $V_{i}=(j_{i-1}+1,\ldots,j_{i})$ satisfying $N=\sum_{i=1}^{m}|V_{i}|$.
Associated with each edge $(r,s)\in E$, we define the jump $e_{rs}=\lvert r\rangle\langle s\rvert$ and the Lindbladian:
$\mathcal{L}_{e_{rs}}(X):=2e_{rs}^{*}Xe_{rs}-e_{rs}^{*}e_{rs}X-Xe_{rs}^{*}e_{rs}\,.$
Then, we let $\sigma_{\beta}:=\sum_{r=1}^{N}\mu_{r}\lvert r\rangle\langle r\rvert$ be a full-rank state with
(5.18) $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{N}>0\,,$
and define $e^{\beta_{rs}}=\mu_{r}/\mu_{s}$ for $1\leq r,s\leq N$ and
$\mathcal{L}_{rs}(X)=e^{\beta_{rs}/2}\mathcal{L}_{e_{rs}}(X)+e^{-\beta_{rs}/2}\mathcal{L}_{e_{sr}}(X)\,.$
Given a symmetric weight function $w(r,s)=w(s,r)>0$ over edges $(r,s)\in E$, we define the $\sigma_{\beta}$-GNS detailed balanced Lindbladian by [^3] In what follows, the sum over edges $E$ counts both $(r,s)\in E$ and $(s,r)\in E$.
(5.19) $\mathcal{L}^{D}=\sum_{(r,s)\in E}w(r,s)\mathcal{L}_{rs}\,,$
noting that each jump $e_{rs}$ is an eigenvector of the modular operator: $\Delta_{\sigma_{\beta}}e_{rs}=e^{\beta_{rs}}e_{rs}$, and its adjoint $e_{sr}$ is also a jump. It follows that $\sigma_{\beta}$ is an invariant state of $\mathcal{L}^{D}$.
Now, a direct computation gives for $j\in V$,
(5.20) $\mathcal{L}^{D}(e_{jj})=\sum_{(r,s)\in E}w(r,s)\mathcal{L}_{rs}(e_{jj})=4\sum_{(r,j)\in E}w(r,j)\left(e^{-\beta_{rj}/2}e_{rr}-e^{\beta_{rj}/2}e_{jj}\right),$
and for $(j,k)\in E$,
(5.21) $\mathcal{L}^{D}(e_{jk})=\kappa_{jk}e_{jk}\quad\text{with}\quad\kappa_{jk}=-2\left(\sum_{(r,j)\in E}w(r,j)e^{\beta_{rj}/2}+\sum_{(r,k)\in E}w(r,k)e^{\beta_{rk}/2}\right).$
We define $\mathcal{B}_{\rm diag}(\mathcal{H}):=\{X\in\mathcal{B}(\mathcal{H})\,;\ X=\sum_{j}x_{j}\lvert j\rangle\langle j\rvert\}$ by the subspace of diagonal matrices with respect to the basis $\lvert j\rangle$ from the invariant state $\sigma_{\beta}$, and $\mathcal{E}_{\rm diag}:\mathcal{B}(\mathcal{H})\to\mathcal{B}_{\rm diag}(\mathcal{H})$ by the projection to the diagonal part of an operator $X\in\mathcal{B}(\mathcal{H})$. From (5.20), we see that the QMS $\mathcal{P}_{t}=e^{t\mathcal{L}^{D}}$ maps $\mathcal{B}_{\rm diag}(\mathcal{H})$ to $\mathcal{B}_{\rm diag}(\mathcal{H})$ and its restriction $\mathcal{P}_{t}\mathcal{E}_{\rm diag}$ is a classical reversible random walk generated by a weighted graph Laplacian: for $(i,j)\in E$,
(5.22) $\mathcal{L}_{\rm cl}(i,j)=\begin{dcases}4w(i,j)e^{-\beta_{ij}/2}\,,&(i,j)\in E\,,\\ -4\sum_{\text{fixed $i$},\,(i,r)\in E}w(i,r)e^{-\beta_{ir}/2}\,,&i=j\in V\,,\\ 0\,,&\text{otherwise}\,,\\ \end{dcases}$
while Eq. 5.21 shows that $e_{jk}$ is an eigenvector of $\mathcal{L}^{D}$. It follows that
(5.23) $\lambda_{D}=\min\big{\{}\lambda_{\min}(-\mathcal{L}_{\rm cl})\,,\min_{i\neq j}-\kappa_{ij}\big{\}}\,,\quad\lVert\mathcal{L}^{D}\rVert_{2\to 2}=\max\big{\{}\lambda_{\max}(-\mathcal{L}_{\rm cl})\,,\max_{i\neq j}-\kappa_{ij}\big{\}}\,,$
where $\lambda_{\min/\max}(-\mathcal{L}_{\rm cl})$ denotes the min/max eigenvalue of $-\mathcal{L}_{\rm cl}$.
By Lemma 2.2, we have
(5.24) $\ker(\mathcal{L}^{D})=\{X\in\mathcal{B}(\mathcal{H})\,;[X,e_{rs}]=[X,e_{sr}]=0\ \text{for $(r,s)\in E$}\}\,.$
We will see that $\dim\ker(\mathcal{L}^{D})>1$, which is a consequence of the disconnectedness of the graph $G$. Recalling the connected components (5.17), for any two nodes $r,s$ in the same component $V_{i}$, there exists a chain of edges $(r,r_{1}),(r_{1},r_{2}),\ldots,(r_{k},v)$ from $E_{i}$ such that
$e_{rv}=e_{rr_{1}}e_{r_{1}r_{2}}\ldots e_{r_{k}v}\,,$
which yields that the $\mathbb{C}$-algebra of $\{e_{rs},e_{sr}\,;\ (r,s)\in E\}$ is given by the block diagonal matrices with respect to the partition $\{V_{i}\}_{i=1}^{m}$ (5.17). Then, by Eq. 5.24, $X\in\ker(\mathcal{L}^{D})$ restricted on each $V_{i}\times V_{i}$ is a multiple of identity. It means that
(5.25) $\ker(\mathcal{L}^{D})=\Big{\{}X\in\mathcal{B}(\mathcal{H})\,;X=\sum_{i=1}^{m}c_{i}\sum_{r\in V_{i}}\lvert r\rangle\langle r\rvert\,,\ c_{i}\in\mathbb{C}\Big{\}}\,,$
and hence the QMS is non-primitive (i.e., $\dim\ker(\mathcal{L}^{D})>1$).
We proceed to characterize the Hamiltonian $H$ such that $[H,\sigma_{\beta}]=0$ and $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ is primitive with $\sigma_{\beta}$ being invariant.
Lemma 5.3.
Under the above setting, it holds that
  • If $\sigma_{\beta}$ has $N$ distinct eigenvalues, i.e., $\mu_{1}>\cdots>\mu_{N}$ in (5.18), then for any Hamiltonian $H$ satisfying $[H,\sigma_{\beta}]=0$, the Lindbladian $\mathcal{L}_{\alpha}$ is non-primitive.
  • In the case of $\sigma_{\beta}$ having degenerate eigenvalues, assume that $\widetilde{\mu}_{i}$ are its distinct eigenvalues with eigenvectors $\lvert r\rangle\in\widetilde{V}_{i}\subset V$, where $\{\widetilde{V}_{i}\}$ is a disjoint partition of $V$. Then, a Hamiltonian $H$ satisfies $[H,\sigma_{\beta}]=0$ and that $\mathcal{L}_{\alpha}$ is primitive if and only if for any $1\leq i,j\leq m$, there exists a chain $(i_{0}=i,i_{1}),(i_{1},i_{2}),\ldots(i_{k-1},i_{k}=j)$ and $\{\widetilde{V}_{j_{l}}\}_{l=1}^{k}$ such that $H_{ab}\neq 0$ for some
    (5.26) $(a,b)\in\Big{(}V_{i_{l-1}}\times V_{i_{l}}\Big{)}\bigcap\left(\widetilde{V}_{j_{l}}\times\widetilde{V}_{j_{l}}\right)\neq\emptyset\,.$
Proof.
Without loss of generality, we let $\alpha=1$. For the first statement, if all eigenvalues of $\sigma_{\beta}$ are distinct, then $H$ satisfying $[H,\sigma_{\beta}]$ has to be a diagonal matrix with respect to $\lvert r\rangle$, and any $X\in\ker(\mathcal{L}^{D})$ satisfies $[H,X]=0$. It follows from Lemma 2.3 that $\ker(\mathcal{L}_{\alpha})=\ker(\mathcal{L}^{D})$ and thus $\mathcal{L}_{\alpha}$ is non-primitive. For the second statement, if $\sigma_{\beta}$ has degenerate eigenvalues, any $H$ satisfying $[H,\sigma_{\beta}]$ is a block diagonal matrix for the partition $\widetilde{V}_{i}$. Then $\mathcal{L}_{\alpha}$ is primitive if and only if $X\in\ker(\mathcal{L}^{D})$ such that $[H,X]=0$ is a multiple of identity. We consider $H=\sum_{rs}H_{rs}\lvert r\rangle\langle s\rvert$ with $\{H_{rs}\}$ being block diagonal and $X\in\ker(\mathcal{L}^{D})$ of the form
(5.27) $X=\sum_{r}x_{r}\lvert r\rangle\langle r\rvert\quad\text{with}\quad(x_{1},\ldots,x_{N})=(\vec{c}_{1},\ldots,\vec{c}_{m})\,,\quad\vec{c}_{i}:=(c_{i},\ldots,c_{i})\in\mathbb{C}^{|V_{i}|}\,.$
Similarly to Eq. 5.9, we need to find $H$ such that $(x_{r}-x_{s})H_{rs}=0$ can imply $x_{1}=\ldots=x_{N}$, more precisely, $c_{1}=\ldots=c_{m}$. It follows that for any distinct $1\leq i,j\leq m$, there exists a sequence $(i_{0}=i,i_{1},\ldots,i_{k}=j)$ such that for each $(i_{l-1},i_{l})$, we can find an edge
$(a,b)\in\Big{(}V_{i_{l-1}}\times V_{i_{l}}\Big{)}\bigcap\left(\widetilde{V}_{j_{l}}\times\widetilde{V}_{j_{l}}\right),$
for some $\widetilde{V}_{j_{l}}$ with $H_{ab}\neq 0$. The proof is complete. ∎
We now assume that $\sigma_{\beta}$ has degenerate eigenvalues such that $(V_{i_{l-1}}\times V_{i_{l}})\bigcap(\widetilde{V}_{j_{l}}\times\widetilde{V}_{j_{l}})\neq\emptyset$, and fix a Hamiltonian $H=\sum_{r,s=1}^{N}H_{rs}\lvert r\rangle\langle s\rvert$ such that $[H,\sigma_{\beta}]=0$ and $\mathcal{L}_{\alpha}$ is primitive (the existence of $H$ is guaranteed by Lemma 5.3). Then $[H,\sigma_{\beta}^{k}]=0$ holds for any $k\in\mathbb{R}$, equivalently,
(5.28) $H_{rs}\mu_{s}^{k}=\mu_{r}^{k}H_{rs}\,.$
One can easily compute for $X=\sum_{r}x_{r}\lvert r\rangle\langle r\rvert\in\ker(\mathcal{L}^{D})$,
$\sigma^{1/2}(HX-XH)\sigma^{1/2}=\sum_{rs}(\mu_{r}\mu_{s})^{1/2}(x_{s}-x_{r})H_{rs}\lvert r\rangle\langle s\rvert\,,$
which implies $\langle X,[H,X]\rangle_{\sigma,1/2}=0$, i.e., $\Pi_{0}\mathcal{L}^{H}\Pi_{0}=0$. Similarly, we have
$\left\lVert\mathcal{L}^{H}\Pi_{+}\right\rVert_{\mathfrak{H}\to\mathfrak{H}}\leq\lambda_{\max}(H)-\lambda_{\min}(H)\,.$
We next estimate the singular value gap of $\mathcal{L}^{H}\Pi_{0}$:
(5.29) $s_{H}^{2}=\inf_{X\in\mathfrak{H}_{0}}\frac{\lVert\mathcal{L}^{H}\Pi_{0}X\rVert_{\sigma,1/2}^{2}}{\lVert X\rVert_{\sigma,1/2}^{2}}\,.$
With the notion in (5.27), the subspace $\mathfrak{H}_{0}$ in (5.29) is given by those $X=\sum_{r}x_{r}\lvert r\rangle\langle r\rvert$ with
$\sum_{r=1}^{N}\mu_{r}x_{r}=\sum_{i=1}^{m}\sum_{r\in V_{i}}\mu_{r}c_{i}=\sum_{i=1}^{m}\widehat{\mu}_{i}c_{i}=0\,,$
where $\{\widehat{\mu}_{i}:=\sum_{r\in V_{i}}\mu_{r}\}$ is a probability distribution over indices $\{i\}$. A direct computation gives
(5.30) $\lVert X\rVert_{\sigma,1/2}^{2}=\sum_{r}\mu_{r}x_{r}^{2}=\sum_{i}\widehat{\mu}_{i}c_{i}^{2}=\langle c,c\rangle_{2,\widehat{\mu}}\,,$
and by Eq. 5.28,
(5.31) $(\mu_{r}\mu_{s})^{1/2}|H_{rs}|^{2}=\mu_{r}|H_{rs}|^{2}=\mu_{s}|H_{rs}|^{2}\,,$
and thus
$\lVert\mathcal{L}^{H}\Pi_{0}X\rVert_{\sigma,1/2}^{2}$ $=\big{\langle}[H,X],\sigma^{1/2}[H,X]\sigma^{1/2}\big{\rangle}=\sum_{r\neq s}\mu_{r}|x_{s}-x_{r}|^{2}|H_{rs}|^{2}$
$=\sum_{i\neq j}|c_{j}-c_{i}|^{2}\widehat{\mu}_{i}\widehat{H}_{ij}\,,$
where
(5.32) $\widehat{H}_{ij}:=\begin{dcases}\frac{1}{\widehat{\mu}_{i}}\sum_{r\in V_{i}}\sum_{s\in V_{j}}\mu_{r}|H_{rs}|^{2}\,,&i\neq j\,,\\ -\sum_{j\ \text{with $j\neq i$}}\widehat{H}_{ij}\,,&i=j\,.\end{dcases}$
By construction, $\widehat{H}_{ij}$ has non-negative entries for $i\neq j$ with each row summing to zero and satisfies the detailed balance with respect to the probability $\widehat{\mu}$ by (5.31):
$\widehat{\mu}_{i}\widehat{H}_{ij}=\sum_{r\in V_{i}}\sum_{s\in V_{j}}\mu_{r}|H_{rs}|^{2}=\sum_{s\in V_{j}}\sum_{r\in V_{i}}\mu_{s}|H_{sr}|^{2}=\widehat{\mu}_{j}\widehat{H}_{ji}\,.$
It follows [norris1998markov] that we can construct a reversible continuous-time Markov chain with $\widehat{H}$ being the generator and $\widehat{\mu}$ being an invariant measure. In addition, by [norris1998markov]*Theorem 3.2.1, this Markov chain is irreducible if and only if $\widehat{H}$ is an irreducible matrix, that is, for any $1\leq i,j\leq m$, there exists a chain $i_{0}=i,i_{1},\ldots,i_{k-1},i_{k}=j$ such that $\widehat{H}_{i_{l-1}i_{l}}\neq 0$, which is guaranteed by Eqs. 5.26 and 5.32. Thus, recalling (5.29), we can write
(5.33) $s_{H}^{2}=\inf_{c\in\mathbb{R}^{m}\,,\sum_{i}\widehat{\mu}_{i}c_{i}=0}\frac{\sum_{i\neq j}|c_{j}-c_{i}|^{2}\widehat{\mu}_{i}\widehat{H}_{ij}}{\langle c,c\rangle_{2,\widehat{\mu}}}=\inf_{c\in\mathbb{R}^{m}\,,\sum_{i}\widehat{\mu}_{i}c_{i}=0}\frac{-2\langle c,\widehat{H}c\rangle_{2,\widehat{\mu}}}{\langle c,c\rangle_{2,\widehat{\mu}}}\,,$
which can be similarly estimated by the canonical path method as in Section 5.1. We briefly repeat the argument below for completeness. For a collection of canonical paths $\Gamma=\{\gamma_{ij}\}_{1\leq i\neq j\leq m}$, we define
$K_{\widehat{\mu},\widehat{H},\Gamma}:=\max_{e}\sum_{e\in\gamma_{ij}}\big{(}\widehat{\mu}_{e^{+}}\widehat{H}_{e^{+}e^{-}}\big{)}^{-1}\widehat{\mu}_{i}\widehat{\mu}_{j}|\gamma_{ij}|\,.$
For any $c\in\mathbb{R}^{m}$ with $\sum_{i}\widehat{\mu}_{i}c_{i}=0$ and a collection of paths $\Gamma=\{\gamma_{ij}\}_{i\neq j}$, we have
$2\sum_{i}\widehat{\mu}_{i}c_{i}^{2}=\sum_{i\neq j}\widehat{\mu}_{i}\widehat{\mu}_{j}(c_{i}-c_{j})^{2}$ $=\sum_{i\neq j}\widehat{\mu}_{i}\widehat{\mu}_{j}\Big{(}\sum_{e=(e^{+},e^{-})\in\gamma_{ij}}(c_{e^{+}}-c_{e^{-}})\Big{)}^{2}$
$\leq\sum_{i\neq j}\widehat{\mu}_{i}\widehat{\mu}_{j}|\gamma_{ij}|\sum_{e=(e^{+},e^{-})\in\gamma_{ij}}(c_{e^{+}}-c_{e^{-}})^{2}$
$=\sum_{e}(c_{e^{+}}-c_{e^{-}})^{2}\widehat{\mu}_{e^{+}}\widehat{H}_{e^{+}e^{-}}\sum_{e\in\gamma_{ij}}\big{(}\widehat{\mu}_{e^{+}}\widehat{H}_{e^{+}e^{-}}\big{)}^{-1}\widehat{\mu}_{i}\widehat{\mu}_{j}|\gamma_{ij}|\,,$
and then
(5.34) $s_{H}\geq\sqrt{2/K_{\widehat{\mu},\widehat{H},\Gamma}}\,.$
We summarize the above discussion with Corollary 3.3 into the following proposition.
Proposition 5.4.
Let $\mathcal{L}_{\alpha}=\alpha\mathcal{L}^{H}+\mathcal{L}^{D}$ be the primitive Lindbladian constructed as above with $\sigma_{\beta}$ being the unique invariant state. Then, the exponential decay holds for $X_{t}=e^{t\mathcal{L}_{\alpha}}X_{0}$ with $X_{0}\in\mathcal{B}(\mathcal{H})$,
$\lVert X_{t}-\langle X_{t}\rangle_{\sigma_{\beta}}\rVert_{\sigma_{\beta},1/2}^{2}\leq Ce^{-\nu t}\lVert X_{0}-\langle X_{0}\rangle_{\sigma_{\beta}}\rVert_{\sigma_{\beta},1/2}^{2}\,,\quad\forall\,t\geq 0\,,$
with $C=\mathcal{O}(1)$ and convergence rate:
$\nu=\frac{\lambda_{D}s_{H}^{2}}{(28s_{H}+5\sqrt{2}\bigl{(}\lambda_{\max}(H)-\lambda_{\min}(H)\bigr{)})^{2}+72\alpha^{-2}\lambda_{D}\big{\|}\mathcal{L}^{D}\big{\|}_{2\to 2}}\,,$
where $\lambda_{D}$ and $\lVert\mathcal{L}^{D}\rVert_{2\to 2}$ are characterized by (5.23), and $s_{H}$ is given by (5.33) with lower bound (5.34).
Example 5.1.
Recalling the decomposition (5.17), suppose that the edges in $E_{i}$ only connect successive vertices: $E_{i}=\{(k,k+1)\,;\ j_{i-1}+1\leq k\leq j_{i}-1\}$. We consider the simplest uniform weight $w(r,s)=1$ for $(r,s)\in E$ and let $\sigma\propto\sum_{r=1}^{N}e^{-\beta r}\lvert r\rangle\langle r\rvert$ ($\beta>0$). Then, for this example, $\kappa_{jk}$ in (5.21) satisfies
$-\kappa_{jk}\in\left\{8\cosh\left(\frac{\beta}{2}\right),4\cosh\left(\frac{\beta}{2}\right),4e^{\frac{\beta}{2}}+2e^{-\frac{\beta}{2}},2e^{\frac{\beta}{2}}+4e^{-\frac{\beta}{2}}\right\}\,,$
which gives $4\cosh(\beta/2)\leq-\kappa_{jk}\leq 8\cosh(\beta/2)$. Moreover, the generator (5.22) on the diagonals is a block diagonal matrix for the partition $\{V_{i}\}$ with each block of the form:
$\mathcal{L}_{{\rm cl},V_{i}}=4\begin{bmatrix}-e^{-\frac{\beta}{2}}&e^{-\frac{\beta}{2}}&0&\cdots&0\\ e^{\frac{\beta}{2}}&-e^{-\frac{\beta}{2}}-e^{\frac{\beta}{2}}&e^{-\frac{\beta}{2}}&\cdots&0\\ 0&e^{\frac{\beta}{2}}&-e^{-\frac{\beta}{2}}-e^{\frac{\beta}{2}}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&-e^{\frac{\beta}{2}}\end{bmatrix}_{|V_{i}|\times|V_{i}|}\,,$
which is nothing else than the generator of the classical birth-death process with constant birth and death rates, and can be viewed as a perturbed tridiagonal Toeplitz matrix:
$\mathcal{L}_{{\rm cl},V_{i}}={\rm tridiag}\left(e^{-\frac{\beta}{2}},-e^{-\frac{\beta}{2}}-e^{\frac{\beta}{2}},e^{\frac{\beta}{2}}\right)+e^{\frac{\beta}{2}}\lvert 1\rangle\langle 1\rvert+e^{-\frac{\beta}{2}}\lvert|V_{i}|\rangle\langle|V_{i}|\rvert\,.$
Its eigenvalues can be explicitly calculated as follows [yueh2008explicit]: $\mu_{1}=0$ and for $2\leq k\leq|V_{i}|$:
$\mu_{k}=-8\cosh\left(\frac{\beta}{2}\right)+8\cos\left(\frac{\pi}{|V_{i}|}(k-1)\right)$
Thanks to (5.23), we have
$\lambda_{D}=4\min\left\{2\cosh\left(\frac{\beta}{2}\right)-2\cos\left(\frac{\pi}{\max_{i}|V_{i}|}\right),\cosh\left(\frac{\beta}{2}\right)\right\},$
and
$\lVert\mathcal{L}^{D}\rVert_{2\to 2}=8\cosh\left(\frac{\beta}{2}\right)-8\cos\left(\pi\frac{\max_{i}|V_{i}|-1}{\max_{i}|V_{i}|}\right).$
This allows us to give a more concrete convergence rate estimate. We remark that the spectral gap lower bounds for a general GNS-detailed balanced Lindbladian can be found in [temme2013lower].
We finally consider the case where $\sigma_{\beta}$ has all distinct eigenvalues $\mu_{1}>\cdots>\mu_{N}>0$. In this case, by Lemma 5.3, it is impossible to find a Hamiltonian $H$ such that $\mathcal{L}^{H}+\mathcal{L}^{D}$ is primitive. However, we shall see that the necessary degeneracy of eigenvalues of $\sigma_{\beta}$ can be easily enforced by lifting $\mathcal{L}^{D}$ to an extended product space, and the desired Hamiltonian could be constructed similarly.
To illustrate the ideas, let us consider a connected graph $G$ instead, implying that the associated Lindbladian $\mathcal{L}^{D}$ in (5.19) is primitive. We introduce the extended invariant state:
$\widetilde{\sigma}_{\beta}=\frac{1}{2}\sigma_{\beta}\otimes{\bf 1}_{2}\,,\quad{\bf 1}_{2}=\left[\begin{matrix}1&0\\ 0&1\end{matrix}\right].$
Let $A$ be a self-adjoint operator on $\mathbb{C}^{2}$ with nonzero eigenvalues $a\neq b\in\mathbb{R}$:
$A=a\lvert 0\rangle\langle 0\rvert+b\lvert 1\rangle\langle 1\rvert\,.$
We define the lifted $\mathcal{L}^{D}$ by
~$\widetilde{\mathcal{L}}^{D}:=\sum_{(r,s)\in E}w(r,s)\widetilde{\mathcal{L}}_{rs}\,,$
with
~$\widetilde{\mathcal{L}}_{rs}(X):=e^{\beta_{rs}/2}\mathcal{L}_{e_{rs}\otimes A}(X)+e^{-\beta_{rs}/2}\mathcal{L}_{e_{sr}\otimes A}(X)\,.$
From the construction, there holds $\Delta_{\widetilde{\sigma}_{\beta}}(e_{rs}\otimes A)=e^{\beta_{rs}}e_{rs}\otimes A$, and ~$\widetilde{\mathcal{L}}^{D}$ satisfies $\widetilde{\sigma}_{\beta}$-GNS DBC. Similarly, the kernel of ~$\widetilde{\mathcal{L}}^{D}$ is characterized by
$\ker(\widetilde{\mathcal{L}}^{D})$ =$=\{X\in\mathcal{B}(\mathcal{H}\otimes\mathbb{C}^{2})\,;\ [X,e_{rs}\otimes A]=[X,e_{sr}\otimes A]=0\ \text{for $(r,s)\in E$}\}$
=$=\Big{\{}X\in\mathcal{B}(\mathcal{H}\otimes\mathbb{C}^{2})\,;\ X=\sum_{r\in V}\sum_{i=0,1}x_{ri}\lvert ri\rangle\langle ri\rvert\,,\ x_{ri}\equiv c_{i}\in\mathbb{C}\ \,\text{for $r\in V$}\Big{\}}\,,$
that is, $X\in\ker(\widetilde{\mathcal{L}}^{D})$ if and only if under the basis $\{\lvert ri\rangle\langle sj\rvert\}_{r,s\in V\,,i,j=0,1}$,
$X={\bf 1}_{N}\otimes\left[\begin{matrix}c_{0}&0\\ 0&c_{1}\end{matrix}\right]\,,$
where ${\bf 1}_{N}$ denotes the $N\times N$ identity matrix. On the other hand, the condition $[H,\widetilde{\sigma}_{\beta}]=0$ gives that $H$ is of the block diagonal form $H={\rm diag}(H^{1},\ldots,H^{N})$ under the basis $\{\lvert ri\rangle\langle sj\rvert\}$, where $H^{i}$ are self-adjoint $2\times 2$ matrices. It readily follows that for $X\in\ker(\widetilde{\mathcal{L}}^{D})$, $\mathcal{L}^{H}(X)=0$ is equivalent to that for each matrix $H^{i}$, there holds $(c_{0}-c_{1})H^{i}_{01}=0$, where $H_{01}^{i}$ is the $(0,1)$-entry of $H^{i}$. Then, the convergence rate of the Lindblad dynamics $e^{t\mathcal{L}_{\alpha}}$ can be estimated by similar arguments as those for Propositions 5.2 and 5.4. The detailed discussion for the convergence of the lifted Lindbladian is beyond the scope of this work and left to future studies.
Acknowledgment
This work is supported in part by National Science Foundation via award NSF DMS-2309378. We appreciate helpful discussions with Hongrui Chen, Di Fang, Lin Lin, Gabriel Stoltz, Yu Tong, and Lexing Ying.